### v3

# Connectedness of fibers beyond semitoric systems I: the non-degenerate case

Abstract.
In this paper we study the connectedness of the fibers of integrable systems that have only non-degenerate singular points and extend complexity one $T$-spaces with proper moment maps. Our main result states that if there are no tall singular points with a hyperbolic block and connected $T$-stabilizer, then each fiber is connected. Moreover, we prove that the above condition is necessary if either some reduced space is simply connected or the moment map for the integrable system is generic in a natural sense.
2020 Mathematics Subject Classification. Primary 37J35, 53D20. Secondary 37J39, 53D35.
Keywords: Completely integrable Hamiltonian systems, non-degenerate singular points, complexity one actions, connected fibers.
1. Introduction
An important step in studying Hamiltonian group actions is understanding the fibers of the moment map. In this direction, connectedness of the fibers has been established in the following cases:
  1. (1)
    Hamiltonian group actions of compact abelian Lie groups with proper moment maps [At, GS2, LMTW]; and
  2. (2)
    semi-toric systems in dimension four and their natural generalization in dimension six [VN, Wa].
In both cases, the connectedness of the fibers of the moment map is key to understanding the moment map image and plays an important role in classification.
In this paper we study the connectedness of the fibers of integrable systems that are extensions of complexity one torus actions, and have only non-degenerate singular points. Complexity one torus actions are the simplest after the toric case and have been partially classified in [KT1, KT2, KT3]. Non-degenerate singular points of integrable systems are generic, and can be thought of as the symplectic analog of critical points for Morse-Bott functions; they have attracted much interest, both for classical and quantum integrable systems (see [BF, SV] and references therein). In particular, the integrable systems that we study include the semi-toric systems of [VN], and their higher dimensional generalization in [Wa].
In our main result, Theorem 1.6 we give sufficient conditions for fibers of the moment map to be connected, generalizing the results in [VN, Wa]. Morse theory plays a central role in the theorems (1) and (2) above, and our theorem both uses and generalizes the following simple Morse-theoretic result: Let $N$ be a closed, connected 2-dimensional manifold and let $\overline{g}\colon N\to{\mathbb{R}}$ be a Morse function. If $\overline{g}$ has no singular point of index one, then each fiber is connected (see Lemma 7.3). We also explore the extent to which the hypotheses in our main result are necessary (see Theorem 1.10).
Before stating our results precisely, we begin with few definitions from integrable systems.
Definition 1.1.
An integrable system is a quadruple $\left(M,\omega,V,f\right)$, where $\left(M,\omega\right)$ is a connected $2n$-dimensional symplectic manifold, $V$ is an $n$-dimensional real vector space, and $f\colon M\to V^{*}$ is a smooth map such that:
  1. (1)
    the Poisson bracket $\big\{\langle f,\xi_{1}\rangle,\langle f,\xi_{2}\rangle\big\}=0$ for all $\xi_{1},\xi_{2}\in V$, and
  2. (2)
    the function $f$ is regular on a dense subset of $M$.
Here $\langle\cdot,\cdot\rangle\colon V^{*}\times V\to{\mathbb{R}}$ is the natural pairing.
The standard definition of integrable systems takes $V={\mathbb{R}}^{n}$. In this paper, we assume that every integrable system is complete, that is, $f$ is a moment map for a Hamiltonian $V$-action.
Fix an integrable system $\left(M,\omega,V,f\right)$. A point $p\in M$ is regular if $D_{p}f$ has full rank and singular otherwise. Given a point $p\in M$, the $V$-orbit $\mathcal{O}$ through $p$ is isotropic. Hence, the vector space $\left(T_{p}\mathcal{O}\right)^{\omega}/T_{p}\mathcal{O}$ inherits a natural symplectic structure. Moreover, the stabilizer $W\subseteq V$ of $p$ acts by linear symplectomorphisms on $T_{p}M$ and fixes $T_{p}\mathcal{O}.$ We call the induced symplectic action of $W$ on $(T_{p}\mathcal{O})^{\omega}/T_{p}\mathcal{O}$ the linearized action at $p$.
Definition 1.2.
Let $p$ be a point in an integrable system $\left(M,\omega,V,f\right)$. Let $W_{0}\subseteq V$ be the Lie algebra of the stabilizer of $p$. The point $p$ is non-degenerate if there exist identifications $W_{0}\simeq{\mathbb{R}}^{\dim W}$ and $(T_{p}\mathcal{O})^{\omega}/T_{p}\mathcal{O}\simeq{\mathbb{C}}^{\dim W}$ such that the homogeneous moment map for the linearized action at $p$ is a product of the following “blocks”:
  • (elliptic) $z\in{\mathbb{C}}\mapsto\frac{1}{2}|z|^{2}$;
  • (hyperbolic) $z\in{\mathbb{C}}\mapsto{\mathfrak{I}}(z^{2})$;
  • (focus-focus) $(z_{1},z_{2})\in{\mathbb{C}}^{2}\mapsto\left(\frac{1}{2}(|z_{1}|^{2}-|z_{2}|^{2}),{\mathfrak{I}}(z_{1}z_{2})\right)$.
Here ${\mathfrak{I}}(w)$ denotes the imaginary part of $w\in{\mathbb{C}}$ and ${\mathbb{C}}^{\dim W}$ has the standard symplectic structure. Moreover, $p$ has purely elliptic type if $p$ is non-degenerate and all the blocks are elliptic.
Remark 1.3.
The above definition of non-degeneracy is equivalent to the standard one (see, for instance, [SV, Theorem 4.8]). In particular, the focus-focus block in the above definition can be transformed to the one in the integrable systems literature by applying a linear symplectomorphism.
Remark 1.4.
The notion of non-degeneracy can be viewed as a “symplectic Morse-Bott” condition. In particular, in dimension two, a singular point is non-degenerate in the sense of Definition 1.2 if and only if it is non-degenerate in the sense of Morse theory. Moreover, it is hyperbolic if and only if it has index one and elliptic otherwise.
Next we introduce complexity one torus actions. Let $T$ be an (n1)(n-1)-dimensional torus with Lie algebra $\mathfrak{t}$}.
Definition 1.5.
A complexity one $\boldsymbol{T}$-space is a triple $(M,\omega,\Phi)$, where $(M,\omega)$ is a connected $2n$-dimensional symplectic manifold and $\Phi\colon M\to{\mathfrak{t}}^{*}$ is a moment map for an effective $T$-action.
Let $(M,\omega,\Phi)$ be a complexity one $T$-space. Given $\beta\in{\mathfrak{t}}^{*}$, the reduced space at $\beta$ is the quotient $\Phi^{-1}(\beta)/T$; if $\beta=0$, we write M//T{M/\!/T} for the reduced space. A point $p\in M$ is short if $[p]$ is an isolated point in the reduced space $\Phi^{-1}(\Phi(p))/T$ and is tall otherwise. By [KT1, Proposition 6.1], if $\Phi$ is proper each reduced space $\Phi^{-1}(\beta)/T$ containing more than one point is topologically a closed, connected oriented 2-dimensional manifold; moreover, any two non-empty reduced spaces have the same fundamental group [KT1, Lemma 5.7 and Corollary 9.7].
The main objects that we study in this paper are non-degenerate integrable systems that extend complexity one torus actions, i.e., integrable systems $\left(M,\omega,{\mathfrak{t}}\times{\mathbb{R}},f:=(\Phi,g)\right)$ such that every singular point is non-degenerate and such that $(M,\omega,\Phi)$ is a complexity one $T$-space. Our main result is as follows.
Theorem 1.6.
Let $\left(M,\omega,{\mathfrak{t}}\times{\mathbb{R}},f:=(\Phi,g)\right)$ be an integrable system such that $(M,\omega,\Phi)$ is a complexity one $T$-space with a proper moment map. Assume that each tall singular point in $\left(M,\omega,{\mathfrak{t}}\times{\mathbb{R}},f\right)$ is non-degenerate and that no such point has a hyperbolic block and connected $T$-stabilizer. Then the fiber $f^{-1}(\beta,c)$ is connected for each $(\beta,c)\in{\mathfrak{t}}^{*}\times{\mathbb{R}}$; moreover, the reduced space $\Phi^{-1}(\beta)/T$ is simply connected.
Remark 1.7.
Theorem 1.6 implies connectedness of the fibers for the higher dimensional analogs of semi-toric systems that Wacheux studied in [Wa]: Integrable systems $\left(M,\omega,{\mathfrak{t}}\times{\mathbb{R}},f:=\left(\Phi,g\right)\right)$ such that $\left(M,\omega,\Phi\right)$ is a complexity one $T$-space with a proper moment map all of whose singular points are non-degenerate without hyperbolic blocks. In particular, our main result generalizes [Wa, Theorem 4.2.4] to arbitrary dimensions.
Remark 1.8.
Theorem 1.6 is comparable to [PRV, Theorems 1 – 3] in that all these results assume that the singular points of an integrable system are non-degenerate, disallow (certain) singular points with a hyperbolic block, and prove that the fibers are connected. However, the results in [PRV] do not require the first component of the moment map to generate a Hamiltonian circle action, but only hold in dimension four and have additional hypotheses on the moment map.
Unfortunately, Theorem 1.6 is not quite sharp. If a reduced space has positive genus, then $M$ must have non-degenerate singular points with a hyperbolic block and connected $T$-stabilizer; see Remark 1.14. However, as we show below, it is still possible for all the fibers of $f$ to be connected in this case.
Example 1.9.
Let $\omega$ be the standard symplectic form on the torus $N={\mathbb{R}}^{2}/{\mathbb{Z}}^{2}$. The function $g\colon N\to{\mathbb{R}}$ defined by $g([x,y])=\cos(2\pi x)+\cos(2\pi y)$ is Morse and has connected fibers, a unique minimum and maximum, and two points of index one. Hence, $(N,\omega,{\mathbb{R}},g)$ is an integrable system with two hyperbolic singular points. (Since the compact torus $T$ is trivial, all points have connected $T$-stabilizer.) To form higher dimensional examples, take the product of this integrable system with a compact, connected symplectic toric manifold.
Nevertheless, as our next result shows, Theorem 1.6 is quite close to being sharp.
Theorem 1.10.
Let $\left(M,\omega,{\mathfrak{t}}\times{\mathbb{R}},f:=(\Phi,g)\right)$ be an integrable system such that $(M,\omega,\Phi)$ is a complexity one $T$-space with a proper moment map. Assume that each tall singular point in $\left(M,\omega,{\mathfrak{t}}\times{\mathbb{R}},f\right)$ is non-degenerate. Moreover, assume that one of the following holds:
  • The reduced space $\Phi^{-1}(\beta_{0})/T$ is simply connected for some $\beta_{0}\in\Phi(M).$
  • Each fiber of $f$ contains at most one $T$-orbit of non-degenerate singular points with a hyperbolic block.
Then the fiber $f^{-1}(\beta,c)$ is connected for all $(\beta,c)\in{\mathfrak{t}}^{*}\times{\mathbb{R}}$ if and only $M$ has no tall singular points with a hyperbolic block and connected $T$-stabilizer.
Remark 1.11.
If in Theorems 1.6 and 1.10 we additionally assume that $M$ is compact, then by a theorem of Hui Li [Li], the fundamental group of $M$ is equal to that of any reduced space. Hence, in this case, the phrase “the reduced space $\Phi^{-1}(\beta)/T$” can be changed to “MM”.
The fundamental idea behind the proofs of Theorems 1.6 and 1.10 is quite simple: Since $\left(M,\omega,{\mathfrak{t}}\times{\mathbb{R}},f=(\Phi,g)\right)$ is an integrable system, the function $g$ is $T$-invariant, and hence induces a function $\overline{g}\colon\Phi^{-1}(\beta)/T\to{\mathbb{R}}$. By the key technical result in our paper, Proposition 7.1, each non-trivial component of the reduced space $\Phi^{-1}(\beta)/T$ can be given the structure of a smooth oriented surface so that $\overline{g}\colon\Phi^{-1}(\beta)/T\to{\mathbb{R}}$ is a Morse function. Moreover, the critical points of index $1$ correspond exactly to orbits of non-degenerate singular points with a hyperbolic black and connected $T$-stabilizer. Since $\Phi$ is proper, the reduced space $\Phi^{-1}(\beta)/T$ is compact and connected by Theorem 5.5. Therefore, the proofs reduce to standard results in Morse theory; see Lemmas 7.3 and 7.4.
This method can be used more generally on integrable systems that extend a complexity one torus action but have (sufficiently nice) degenerate singular points. For example, in a forthcoming paper, we will prove that the analogs of Theorems 1.6 and 1.10 hold if we also allow a broader class of “ephemeral” singular points; like focus-focus points and hyperbolic points with disconnected $T$-stabilizers, these correspond to regular points on the symplectic quotients.
Remark 1.12.
The second bullet point in Theorem 1.10 can be thought of as a genericity condition because generic Morse functions have at most one critical point (of index 1) in each fiber.
Remark 1.13.
More generally, Theorems 1.6 and 1.10 both hold if $\Phi$ is proper as a map to an open convex subset of $\mathfrak{t}$$^{*}$, because the analog of Theorem 5.5 still holds in this case [LMTW, Theorem 4.3]. Additionally, if $f$ is proper and the fibers of $\Phi$ are connected, then the first statement in Theorem 1.6 still holds. This is because a proper Morse function on a connected surface with no critical points of index $1$ has connected fibers (Proposition 5.6).
Remark 1.14.
Let $(M,\omega,{\mathfrak{t}}^{*}\times{\mathbb{R}},f)$ be an integrable system satisfying the assumptions in the first two sentences of Theorem 1.6, but assume that the reduced space $\Phi^{-1}(\beta)/T$ has positive genus $k$, and hence Euler characteristic 22k2-2k. In this case, by calculating the Euler characteristic, we see that $\Phi^{-1}(\beta)/T$ has at least $2k$ orbits of non-degenerate singular points with a hyperbolic block and connected $T$-stabilizer because because $\overline{g}$ has at least one minimum and maximum. Note that if $\Phi^{-1}(\beta)/T$ has more than $2k$ such orbits, then the same argument shows that $\overline{g}$ must have additional local minima or maxima; hence the fibers of $f$ are not connected by Lemma 7.2.
Structure of the paper
In Section 2, we review some facts about tall complexity one local models and then briefly explore a few additional properties. Next, in Section 3 we prove that we can endow the symplectic quotient $\Phi^{-1}(\beta)/T$ with a smooth structure so that $\overline{g}$ is a Morse function – as long as most singular points satisfy a certain technical condition that guarantees that we can do this locally. Most of the rest of the paper is dedicated to proving that non-degenerate singular points satisfy this condition. For example, in Section 4 we prove that tall points that have purely elliptic type satisfy it, and in Section 5 we show that, if $M$ is four-dimensional, focus-focus points satisfy it. At this point, we are able to demonstrate our technique by using it to give a new proof that that fibers of a semi-toric system are connected. In Section 6, we prove that all the remaining types of non-degenerate singular points also satisfy this technical condition. Therefore, we can complete the proof of our theorems in Section 7.
Acknowledgments
D. Sepe was partially supported by FAPERJ grant JCNE E-26/202.913/2019 and by a CAPES/Alexander von Humboldt Fellowship for Experienced Researchers 88881.512955/2020-01. This study was financed in part by the Coordenação de Aperfeiçoamento de Pessoal de Nível Superior – Brazil (CAPES) – Finance code 001. S. Tolman was partial supported by NSF DMS Award 2204359 and Simons Foundation Collaboration Grant 637995.
2. Tall local models
To begin, we review some facts about tall complexity one local models. In particular, we introduce (tall) complexity one local models, construct the “defining polynomial” for tall local models, and use it to describe the symplectic quotient Y//T{Y/\!/T}. Moreover, we provide a generating set for $T$-invariant polynomials on these models.
Fix a complexity one $T$-space $(M,\omega,\Phi)$ and let $\mathfrak{O}$ be the $T$-orbit through a point $p\in M$. Since $\mathfrak{O}$ is isotropic, the vector space $(T_{p}\mathfrak{O})^{\omega}/T_{p}\mathfrak{O}$ inherits a natural symplectic structure. Moreover, the natural action of the stabilizer $H\subseteq T$ of $p$ on $T_{p}M$ induces a symplectic representation of $H$ on $(T_{p}\mathfrak{O})^{\omega}/T_{p}\mathfrak{O}$, called the symplectic slice representation at $p$. It is isomorphic to the representation of $H$ on ${\mathbb{C}}^{h+1}$ associated to some injective homomorphism $\rho\colon H\to(S^{1})^{h+1}$, where $h=\dim H$. Let $\eta_{i}\in{\mathfrak{h}}^{*}$ be the differential of the $i$’th component of $\rho$. The (not necessarily distinct) vectors $\eta_{0},\dots,\eta_{h}$ are called the isotropy weights at $p$; the multiset of isotropy weights at $p$ is uniquely determined.
Set $Y:=T\times_{H}{\mathfrak{h}}^{\circ}\times{\mathbb{C}}^{h+1}$, where $\mathfrak{h}$$^{\circ}$ is the annihilator of $\mathfrak{h}$} in $\mathfrak{t}$$^{*}$ and $H$ acts on $T$ by multiplication, on $\mathfrak{h}$$^{\circ}$ trivially, and on ${\mathbb{C}}^{h+1}$ by $h\cdot z=\rho(h^{-1})z$. Fix an inner product on the Lie algebra $\mathfrak{t}$} once and for all, and use it to identify $\mathfrak{t}$$^{*}$ with ${\mathfrak{h}}^{\circ}\oplus{\mathfrak{h}}^{*}.$ There exists a canonical symplectic form $\omega_{Y}$ on $Y$ so that the $T$-action $s\cdot[t,\alpha,z]=[st,\alpha,z]$ has homogeneous moment map $\Phi_{Y}\colon Y\to{\mathfrak{t}}^{*}$ given by
$\Phi_{Y}([t,\alpha,z])=\alpha+\Phi_{H}(z),$
where $\Phi_{H}(z)=\frac{1}{2}\sum_{i}\eta_{i}|z_{i}|^{2}$ is the homogeneous moment map for the $H$-action on ${\mathbb{C}}^{h+1}$; see Remark 4.2. By the Marle-Guillemin-Sternberg local normal form theorem, there is a $T$-equivariant symplectomorphism from an invariant neighborhood of $\mathfrak{O}$ to an open subset of $Y$ taking $p$ to $[1,0,0]$. We call $(Y,\omega_{Y},\Phi_{Y})$ the local model for $p$. We say that $Y$ is short if $\Phi_{Y}^{-1}(0)$ consists of a single $T$-orbit and tall otherwise. By Lemma 2.7 below, the fiber $\Phi^{-1}_{Y}(0)$ contains exactly one $T$-orbit exactly if $p]\in\Phi^{-1}(\Phi(p))/T$ is an isolated point. Hence, $p$ is short exactly if its local model is short.
As shown in [KT1], tall local models [^1] Here, and throughout the paper, all the local models that we consider are complexity one. have an important invariant, called the defining polynomial.
Lemma 2.1.
Let $Y=T\times_{H}{\mathfrak{h}}^{\circ}\times{\mathbb{C}}^{h+1}$ be a local model associated to an injective homomorphism $\rho\colon H\to(S^{1})^{h+1}$. Then there exists $\xi=(\xi_{0},\ldots,\xi_{h})\in{\mathbb{Z}}^{h+1}$ such that the homomorphism $(S^{1})^{h+1}\to S^{1}$ given by $\lambda\mapsto\prod\limits_{j=0}^{h}\lambda_{j}^{\xi_{j}}$ induces a short exact sequence
(2.2) $1\to H\stackrel{{\scriptstyle\rho}}{{\to}}(S^{1})^{h+1}\to S^{1}\to 1;$
moreover, the vector $\xi$ is unique up to sign. Finally, $\xi_{i}\xi_{j}\geq 0$ for all $i,j$ if and only if the local model $Y$ is tall.
Proof.
Existence of $\xi$ and its uniqueness up to sign follow from [KT1, Lemma 5.8]. The local model $Y$ is tall if and only if $\Phi^{-1}_{Y}(0)$ consists of more than one orbit. By [KT1, Lemma 5.4], the latter occurs exactly if $\Phi_{Y}$ is not proper, which, by [KT1, Lemma 5.8], is equivalent to $\xi_{i}\xi_{j}\geq 0$ for all $i,j$. ∎
Remark 2.3.
Let $\eta_{0},\ldots,\eta_{h}\in{\mathfrak{h}}^{*}$ be the isotropy weights at $p=[1,0,0]\in Y$. By (2.2), the homomorphism ${\mathbb{R}}^{h+1}\to{\mathbb{R}}$ given by $a\mapsto\sum\limits_{j}a_{j}\xi_{j}$ induces dual short exact sequences
$0\to{\mathfrak{h}}\stackrel{{\scriptstyle\rho_{*}}}{{\to}}{\mathbb{R}}^{h+1}\to{\mathbb{R}}\to 0;$
$0\to{\mathbb{R}}^{*}\to\left({\mathbb{R}}^{h+1}\right)^{*}\stackrel{{\scriptstyle\rho^{*}}}{{\to}}{\mathfrak{h}}^{*}\to 0.$
In particular, $b=(b_{0},\ldots,b_{h})\in{\mathbb{R}}^{h+1}$ lies in $\rho_{*}({\mathfrak{h}})$ exactly if $\sum\limits_{j}b_{j}\xi_{j}=0$, while $\rho^{*}(a)=\sum\limits_{j}a_{j}\eta_{j}$ vanishes exactly if $a$ is a multiple of $\xi$. Hence, if $Y$ is tall then $(1,\dots,1)\notin\rho_{*}({\mathfrak{h}})$ because $\xi\neq 0$ can be chosen to lie in ${\mathbb{Z}}_{\geq 0}^{h+1}.$
Definition 2.4.
Fix a tall local model $Y$. By Lemma 2.1, there exists a unique $\xi=(\xi_{0},\ldots,\xi_{h})\in{\mathbb{Z}}_{\geq 0}^{h+1}$ inducing the short exact sequence (2.2). The map $P\colon{\mathbb{C}}^{h+1}\to{\mathbb{C}}$ given by $P(z)=\prod_{j=0}^{h}z_{j}^{\xi_{j}}$ is called the defining polynomial of $Y$. By a slight abuse of notation, we use the same name for the map $P\colon Y\to{\mathbb{C}}$ given by $t,\alpha,z]\mapsto\prod_{j=0}^{h}z_{j}^{\xi_{j}}$. The sum $N:=\sum_{j=0}^{h}\xi_{j}$ is called the degree of $P$.
Example 2.5.
Let ${\mathbb{Z}}_{2}=\{\pm 1\}$ act on $\mathbb{C}$} by $\lambda\cdot z=\lambda z$. The associated local model $S^{1}\times_{{\mathbb{Z}}_{2}}{\mathbb{R}}^{*}\times{\mathbb{C}}$ is tall. The defining polynomial $P\colon{\mathbb{C}}\to{\mathbb{C}}$ is given by $P(z)=z^{2}$ and has degree $2$.
Example 2.6.
Let $S^{1}$ act effectively on ${\mathbb{C}}^{2}$ by $\lambda\cdot(z_{1},z_{2})=(\lambda z_{1},\lambda^{-1}z_{2})$ with moment map $\Phi(z_{1},z_{2})=\frac{1}{2}(|z_{1}|^{2}-|z_{2}|^{2})$. This is a tall local model. The defining polynomial $P\colon{\mathbb{C}}^{2}\to{\mathbb{C}}$ is given by $P(z_{1},z_{2})=z_{1}z_{2}$ and also has degree $2$.
The reduced space ${Y/\!/T}=\Phi^{-1}_{Y}(0)/T$ is endowed with the subquotient topology and a sheaf of smooth functions defined as follows: For any open subset $U\subseteq{Y/\!/T}$, let $C^{\infty}(U)$ be the set of functions $\overline{g}\colon U\to{\mathbb{R}}$ whose pullback to $\Phi_{Y}^{-1}(0)$ is the restriction of a smooth $T$-invariant function $g\colon Y\to{\mathbb{R}}$. A homeomorphism between topological spaces endowed with sheaves of smooth functions is a diffeomorphism if it induces an isomorphism between the sheaves of smooth functions.
Lemma 2.7.
Let $Y=T\times_{H}{\mathfrak{h}}^{\circ}\times{\mathbb{C}}^{h+1}$ be a tall local model with moment map $\Phi_{Y}\colon Y\to{\mathfrak{t}}^{*}$. The defining polynomial $P\colon Y\to{\mathbb{C}}$ induces a homeomorphism $\overline{P}\colon{Y/\!/T}\to{\mathbb{C}}$ that restricts to a diffeomorphism on the complement of $\left\{[1,0,0]\right\}$.
Proof.
Following [KT1], we say that an orbit is exceptional if every nearby orbit in the same moment fiber has strictly smaller stabilizer. Arguing as in the proof of Lemma 2.1, since $Y$ is tall the moment map $\Phi_{Y}$ is not proper. The first claim follows at once by [KT1, Corollary 6.3]. By [KT1, Lemmas 5.15 and 5.17], the orbit of $[1,0,0]$ is the only possible exceptional orbit in Y//T{Y/\!/T}. The second claim now follows at once by [KT1, Corollary 7.2]. ∎
In most cases, the homeomorphism $\overline{P}\colon{Y/\!/T}\to{\mathbb{C}}$ is not globally a diffeomorphism; however, as the lemma below shows, it is a diffeomorphism when $P(z)=z_{0}$.
Lemma 2.8.
Let $Y=T\times_{H}{\mathfrak{h}}^{\circ}\times{\mathbb{C}}^{h+1}$ be a tall local model with defining polynomial $P(z)=\prod_{i=0}^{h}z_{i}^{\xi_{i}}$. Assume that $\xi_{i}=0$ for all $i>k$. Let the group
$K:=\{\lambda\in(S^{1})^{k+1}\mid\prod_{j=0}^{k}\lambda_{j}^{\xi_{j}}=1\}$
act naturally on ${\mathbb{C}}^{k+1}$. The inclusion ${\mathbb{C}}^{k+1}\hookrightarrow Y$ given by
$(z_{0},\ldots,z_{k})\mapsto[1,0,(z_{0},\ldots,z_{k},0,\ldots,0)]$
induces a diffeomorphism between ${\mathbb{C}}^{k+1}/\!/K$ and Y//T{Y/\!/T}. In particular, if $P(z)=z_{0}$ then $\overline{P}\colon{Y/\!/T}\to{\mathbb{C}}$ is a diffeomorphism with inverse $z\mapsto[1,0,(z,0)].$
Proof.
Let $H$ act on ${\mathbb{C}}^{h+1}$ via $\rho\colon H\hookrightarrow(S^{1})^{h+1}$. Let $\Phi_{H}\colon{\mathbb{C}}^{h+1}\to{\mathfrak{h}}^{*}$ and $\Phi_{K}\colon{\mathbb{C}}^{k+1}\to\mathfrak{k}^{*}$ be the homogeneous moment maps for the $H$-action on ${\mathbb{C}}^{h+1}$ and the $K$-action on ${\mathbb{C}}^{k+1}$, respectively. Since $\xi_{i}=0$ for all $i>k$, we have $\Phi_{H}^{-1}(0)=\Phi_{K}^{-1}(0)\times\{0\}$, and so $\Phi_{Y}^{-1}(0)=T\times_{H}\{0\}\times(\Phi_{K}^{-1}(0)\times\{0\})$, where $\Phi_{Y}:Y\to{\mathfrak{t}}^{*}$ is the moment map for the $T$-action on $Y$. Since $\rho(H)=K\times(S^{1})^{h-k}$, the claim follows immediately. ∎
On the reduced space Y//T{Y/\!/T} there is a simple relation between the magnitudes of the defining polynomial and of $z$.
Lemma 2.9.
Let $Y:=T\times_{H}{\mathfrak{h}}^{\circ}\times{\mathbb{C}}^{h+1}$ be a tall local model with moment map $\Phi_{Y}$. Let $N$ be the degree of the defining polynomial $P(z)=\prod_{j}z_{j}^{\xi_{j}}$. Then for all $t,\alpha,z]\in{Y/\!/T}$,
$\big\lvert z\rvert^{2}=N\big(\prod_{j}\xi_{j}^{\xi_{j}}\big)^{-\frac{1}{N}}\lvert P(z)\big\rvert^{\frac{2}{N}},$
where $0^{0}=1$ by convention.
Proof.
Let $\eta_{0},\dots,\eta_{h}\in\mathfrak{h}^{*}$ be the weights for the $H$-action on ${\mathbb{C}}^{h+1}$. The moment map $\Phi_{Y}$ sends $t,\alpha,z]\in Y$ to $\alpha+\Phi_{H}(z)\in{\mathfrak{t}}^{*}$, where $\Phi_{H}(z)=\sum_{j=0}^{h}\frac{1}{2}|z_{j}|^{2}\eta_{j}$. Thus it suffices to check that $\lvert z\rvert^{2}=\big(\prod_{j}\xi_{j}^{\xi_{j}}\big)^{-\frac{1}{N}}\big\lvert P(z)\big\rvert^{\frac{2}{N}}$ for all $z\in\Phi_{H}^{-1}(0)$. Fix $z\in\Phi_{H}^{-1}(0)$. Since the local model $Y$ is tall and $\Phi_{H}(z)=\sum_{j=0}^{h}\frac{1}{2}|z_{j}|^{2}\eta_{j}=0$, by Remark 2.3 there exists $c\geq 0$ such that $|z_{j}|^{2}=c\xi_{j}$ for all $j$. Therefore,
$\begin{split}\lvert z\rvert^{2}&=\sum\limits_{j=0}^{h}\lvert z_{j}\rvert^{2}=cN;\\ |P(z)|^{2}=\Big\lvert\prod_{j=0}^{h}z_{j}^{\xi_{j}}\Big\rvert^{2}&=\prod\limits_{j=0}^{h}\lvert z_{j}\rvert^{2\xi_{j}}=c^{N}\prod\limits_{j=0}^{h}\xi_{j}^{\xi_{j}}.\end{split}$
The result holds by combining these equations. ∎
To study integrable systems that extend complexity one $T$-actions, we need to understand $T$-invariant functions on these spaces. We start by considering the simplest $T$-invariant functions on local models: $T$-invariant polynomials.
Definition 2.10.
Let $Y=T\times_{H}{\mathfrak{h}}^{\circ}\times{\mathbb{C}}^{h+1}$ be a local model. By a slight abuse of notation, we call a $T$-invariant function $g\colon Y\to{\mathbb{R}}$ a $\boldsymbol{T}$-invariant (homogeneous) polynomial if its restriction to ${\mathfrak{h}}^{\circ}\times{\mathbb{C}}^{h+1}$ is a (homogeneous) polynomial.
Lemma 2.11.
Let $Y=T\times_{H}{\mathfrak{h}}^{\circ}\times{\mathbb{C}}^{h+1}$ be a tall local model with moment map $\Phi_{Y}\colon Y\to{\mathfrak{t}}^{*}$. The algebra of $T$-invariant polynomials on $Y$ is generated by the real and imaginary parts of the defining polynomial $P\colon Y\to{\mathbb{C}}$, the map $t,\alpha,z]\mapsto|z|^{2}$, and the components of $\Phi_{Y}$.
Proof.
By definition, there is a one-to-one correspondence between $T$-invariant polynomials on $Y$ and $H$-invariant polynomials on ${\mathfrak{h}}^{\circ}\times{\mathbb{C}}^{h+1}$. Let $H$ act on ${\mathbb{C}}^{h+1}$ via $\rho\colon H\hookrightarrow(S^{1})^{h+1}$ and let $\Phi_{H}\colon{\mathbb{C}}^{h+1}\to{\mathfrak{h}}^{*}$ be the homogeneous moment map. Since $\Phi_{Y}([t,\alpha,z])=\alpha+\Phi_{H}(z)$ for all $t,\alpha,z]\in Y$, it suffices to prove that the algebra of $H$-invariant polynomials on ${\mathbb{C}}^{h+1}$ is generated by the real and imaginary parts of the defining polynomial $P\colon{\mathbb{C}}^{h+1}\to{\mathbb{C}}$, the map $z\mapsto|z|^{2}$, and the components of $\Phi_{H}$. Since the local model $Y$ is tall, Remark 2.3 implies that $(1,\ldots,1)\notin\rho_{*}({\mathfrak{h}})$. Hence, since $\Phi_{H}(z)=\frac{1}{2}\rho^{*}(|z_{0}|^{2},\ldots,|z_{h}|^{2})$, the subalgebra of $H$-invariant polynomials on ${\mathbb{C}}^{h+1}$ generated by the map $z\mapsto|z|^{2}$ and the components of $\Phi_{H}$ is also generated by the functions $z\mapsto|z_{i}|^{2}$ for $i=0,\ldots,h$.
Let $P\colon{\mathbb{C}}^{h+1}\to{\mathbb{C}}$ be given by $P(z)=\prod_{i}z_{i}^{\xi_{i}}$. Consider a complex-valued polynomial of the form $Q(z)=\prod_{i}z_{i}^{j_{i}}\,\overline{z}_{i}^{k_{i}}$. Given $\lambda\in(S^{1})^{h+1}$, $Q(\lambda z)=\prod_{i}\lambda_{i}^{j_{i}-k_{i}}Q(z)$. In particular, $Q$ is $H$-invariant exactly if $\prod_{i}\lambda_{i}^{j_{i}-k_{i}}=1$ for all $\lambda\in\rho(H)\varsubsetneq(S^{1})^{h+1}$. Hence, by Lemma 2.1, $Q$ is $H$-invariant exactly if the vector $(j_{0}-k_{0},j_{1}-k_{1},\dots,j_{h}-k_{h})$ is a multiple of $(\xi_{0},\dots,\xi_{h})$, i.e., exactly if $Q$ is some product of $P$, its conjugate, and polynomials of the form $z\mapsto|z_{i}|^{2}$. Since the the vector space of complex-valued polynomials on ${\mathbb{C}}^{h+1}$ decomposes as the direct sum of weight spaces that are spanned by elements of the form $z\mapsto\prod_{i}z_{i}^{j_{i}}\overline{z}_{i}^{k_{i}}$, it follows that $H$-invariant complex-valued polynomials are generated by $P$, its conjugate, and the polynomials $z\mapsto|z_{i}|^{2}$. The claim follows immediately. ∎
Remark 2.12.
The proof of Lemma 2.11 provides an equivalent generating set for the algebra of $T$-invariant polynomials on a tall local model $Y$, namely the real and imaginary parts of the defining polynomial $P\colon Y\to{\mathbb{C}}$, the maps $|z_{i}|^{2}\colon Y\to{\mathbb{R}}$ that send $t,\alpha,z]$ to $|z_{i}|^{2}$ for $i=0,\ldots,h$, and the components of the map $t,\alpha,z]\mapsto\alpha$.
3. The key proposition
In this section, we prove that we can endow each non-trivial symplectic quotient with the structure of a smooth oriented surface so that the induced function $\overline{g}$ is Morse – as long as most singular points satisfy a certain technical condition (see Proposition 3.4). This eventually allows us to use Morse theory in two dimensions to obtain the desired results.
Let $\left(M,\omega,{\mathfrak{t}}\times{\mathbb{R}},f:=\left(\Phi,g\right)\right)$ be an integrable system such that $\left(M,\omega,\Phi\right)$ is a complexity one $T$-space. As in the Introduction, we denote the induced function on the reduced space $\Phi^{-1}(\Phi(p))/T$ by $\overline{g}$. We need the following definition.
Definition 3.1.
A point $p\in M$ is a critical point of $\boldsymbol{g}$ modulo $\boldsymbol{\Phi}$ if $D_{p}g$ vanishes on ${\operatorname{ker}}D_{p}\Phi$; otherwise, $p$ is a regular point of $\boldsymbol{g}$ modulo $\boldsymbol{\Phi}$.
Near each regular point of $g$ modulo $\Phi$ there is a natural structure of smooth manifold on the reduced space so that $\overline{g}$ is smooth with no critical points. More precisely, the following holds.
Lemma 3.2.
Let $\left(M,\omega,{\mathfrak{t}}\times{\mathbb{R}},f:=\left(\Phi,g\right)\right)$ be an integrable system such that $\left(M,\omega,\Phi\right)$ is a complexity one $T$-space. Let $p\in M$ be a regular point of $g$ modulo $\Phi$. Then there exists an open neighborhood $U_{p}$ of $[p]$ in the reduced space $\Phi^{-1}(\Phi(p))/T$ and a map $\Psi_{p}\colon U_{p}\to{\mathbb{R}}^{2}$ satisfying:
  1. (a)
    $\Psi_{p}\colon U_{p}\to\Psi(U_{p})$ is a diffeomorphism onto an open set; and
  2. (b)
    $\overline{g}\circ\Psi_{p}^{-1}\colon\Psi_{p}(U_{p})\to{\mathbb{R}}$ is a smooth function with no critical points.
Moreover, $p$ is tall, has purely elliptic type, and the defining polynomial of the local model for $p$ has degree $1$.
Proof.
The result is local. Hence, by the Marle-Guillemin-Sternberg local normal form theorem, we may assume that $M$ is a local model $Y=T\times_{H}{\mathfrak{h}}^{\circ}\times{\mathbb{C}}^{h+1}$, that $\Phi=\Phi_{Y}$, and that $p=[1,0,0]$. First we show that $Y$ is tall. If it is short, then $\Phi^{-1}(\Phi(p))=T\cdot p$. Since $\Phi$ is invariant under the Hamiltonian flow of $g$, this implies that $(T\times{\mathbb{R}})\cdot p=T\cdot p$. Hence $p$ is a critical point of $g$ modulo $\Phi$, which contradicts our assumption.
We identify $T_{p}Y$ with ${\mathfrak{h}}^{\perp}\times{\mathfrak{h}}^{\circ}\times{\mathbb{C}}^{h+1}$ via the map ${\mathfrak{h}}^{\perp}\times{\mathfrak{h}}^{\circ}\times{\mathbb{C}}^{h+1}\to Y$ given by $(\xi,\alpha,z)\mapsto[{\operatorname{exp}}(\xi),\alpha,z]$. The kernel of $D_{p}\Phi$ is given by ${\mathfrak{h}}^{\perp}\times\{0\}\times{\mathbb{C}}^{h+1}$. By Remark 2.12, if the degree $N$ of the defining polynomial $P$ is greater than 1, then the only $T$-invariant polynomials of degree 1 on $Y$ are the components of the map $t,\alpha,z]\mapsto\alpha$, while if $N=1$, the real and imaginary parts of $P$ are also $T$-invariant polynomials of degree 1. Since the derivative of the map $t,\alpha,z]\mapsto\alpha$ at $p$ vanishes when restricted to ${\operatorname{ker}}D_{p}\Phi$, this implies that $N=1$. If we assume, without loss of generality, that $P(z)=z_{0}$, it also implies that $D_{p}g$ restricts to a non-zero $H$-invariant linear function on $\{0\}\times\{0\}\times({\mathbb{C}}\times\{0\})$. By Lemma 2.8, $\overline{P}$ is a diffeomorphism and $\overline{g}\circ\overline{P}^{-1}\colon{\mathbb{C}}\simeq{\mathbb{R}}^{2}\to{\mathbb{R}}$ is equal to the restriction of $g$ to ${\mathbb{C}}\simeq\{1\}\times_{H}\{0\}\times({\mathbb{C}}\times\{0\})\subseteq Y$. Hence, since the restriction of $D_{p}g$ to $\{0\}\times\{0\}\times({\mathbb{C}}\times\{0\})$ is non-zero, there exists an open neighborhood $U_{p}$ of $[p]$ in Y//T{Y/\!/T} such that the restriction of $\overline{g}\circ\overline{P}^{-1}$ to $\overline{P}(U_{p})$ is smooth and has no critical points. Setting $\Psi_{p}\colon U_{p}\to{\mathbb{R}}^{2}$ to be the restriction of $\overline{P}$ to $U_{p}$, the first result follows.
Let $\mathcal{O}$ be the $(T\times{\mathbb{R}})$-orbit through $p$. Since $D_{p}g$ restricts to a non-zero function on ${\mathbb{C}}\times\{0\}$, the vector space $\left(T_{p}\mathcal{O}\right)^{\omega}/T_{p}\mathcal{O}$ is isomorphic to ${\mathbb{C}}^{h}$ and the Lie algebra of the $(T\times{\mathbb{R}})$-stabilizer of $p$ is ${\mathfrak{h}}\times\{0\}\varsubsetneq{\mathfrak{t}}\times{\mathbb{R}}$. Hence, the linearized action on $\left(T_{p}\mathcal{O}\right)^{\omega}/T_{p}\mathcal{O}$ is isomorphic to the $H$-action on ${\mathbb{C}}^{h}$. Finally, since $P(z)=z_{0}$, $H$ acts on ${\mathbb{C}}^{h+1}\simeq{\mathbb{C}}\times{\mathbb{C}}^{h}$ through an isomorphism $H\simeq\{1\}\times(S^{1})^{h}$. Therefore, $p$ has purely elliptic type. ∎
Remark 3.3.
Clearly, every critical point of $g$ modulo $\Phi$ is a singular point of $(M,\omega,{\mathfrak{t}}\times{\mathbb{R}},f:=(\Phi,g))$. The converse holds over the interior of $\Phi(M)$ if $\Phi$ is proper. To see this, let $p$ be a regular point of $g$ modulo $\Phi$. By Lemma 3.2, $p$ is tall and the degree of the defining polynomial of the local model for $p$ is $1$. Moreover, if $\Phi(p)$ is in the interior of $\Phi(M)$, then the local model for $p$ has a surjective moment map because $\Phi$ is open as a map to $\Phi(M)$ by Theorem 5.5. Together, these facts imply that the $T$-stabilizer of $p$ is trivial. Therefore, $p$ is a regular point of $\Phi$, and so also of $f$.
By Lemma 3.2, to construct the desired Morse function on a reduced space we need to impose local conditions near each critical point of $g$ modulo $\Phi$. More precisely, the following key proposition holds.
Proposition 3.4.
Let $\left(M,\omega,{\mathfrak{t}}\times{\mathbb{R}},f:=\left(\Phi,g\right)\right)$ be an integrable system such that $\left(M,\omega,\Phi\right)$ is a complexity one $T$-space. Fix $\beta\in{\mathfrak{t}}^{*}$. Assume that for each critical point $p\in\Phi^{-1}(\beta)$ of $g$ modulo $\Phi$ there exists an open neighborhood $U_{p}$ of $[p]$ in the reduced space $\Phi^{-1}(\beta)/T$ and a map $\Psi_{p}\colon U_{p}\to{\mathbb{R}}^{2}$ satisfying:
  1. (1)
    $\Psi_{p}\colon U_{p}\to\Psi_{p}(U_{p})$ is a homeomorphism onto an open set;
  2. (2)
    $\Psi_{p}$ restricts to a diffeomorphism on $U_{p}\smallsetminus\{[p]\}$; and
  3. (3)
    $\overline{g}\circ\Psi_{p}^{-1}\colon\Psi_{p}(U_{p})\to{\mathbb{R}}$ is a Morse function.
Then $\Phi^{-1}(\beta)/T$ can be given the structure of a smooth oriented surface such that $\overline{g}\colon\Phi^{-1}(\beta)/T\to{\mathbb{R}}$ is a Morse function. Moreover, $[p]$ is a critical point of $\overline{g}$ of index $\mu$ exactly if both $p$ is a critical point of $g$ modulo $\Phi$ and $\Psi_{p}([p])$ is a critical point of $\overline{g}\circ\Psi_{p}^{-1}$ of index $\mu$.
Proof.
Assume for simplicity that $\beta=0$. By hypothesis and Lemma 3.2, for each $p\in\Phi^{-1}(0)$ there exists an open neighborhood $U_{p}$ of $[p]$ in the reduced space ${M/\!/T}:=\Phi^{-1}(0)/T$ and a map $\Psi_{p}\colon U_{p}\to{\mathbb{R}}^{2}$ satisfying properties (1) – (3) above. Since M//T{M/\!/T} is paracompact, the open cover $\{U_{p}\}_{p\in\Phi^{-1}(0)}$ has a locally finite refinement. Hence, there exists a proper subset $\mathcal{P}\varsubsetneq\Phi^{-1}(0)$ and a locally finite open cover $\{U_{p}\}_{p\in\mathcal{P}}$ such that there exists a map $\Psi_{p}\colon U_{p}\to{\mathbb{R}}^{2}$ satisfying properties (1) – (3) above for each $p\in\mathcal{P}$; however, we may no longer assume that $U_{p}$ contains $p$. Since a locally finite set of points is closed, we may also assume that if $p]\in U_{p}$ then $p]\notin U_{p^{\prime}}$ for all $p^{\prime}\neq p$ in $\mathcal{P}$. By construction, for all $p,p^{\prime}\in\mathcal{P}$, the transition function $\Psi_{p}\circ\Psi_{p^{\prime}}^{-1}$ is a diffeomorphism from $\Psi_{p^{\prime}}(U_{p}\cap U_{p^{\prime}})$ to $\Psi_{p}(U_{p}\cap U_{p^{\prime}})$. Hence, the coordinate charts $\Psi_{p}$ endow the reduced space M//T{M/\!/T} with the structure of a $2$-dimensional smooth manifold. Moreover, with respect to this smooth structure, $\overline{g}\colon{M/\!/T}\to{\mathbb{R}}$ is a Morse function. Additionally, $[p]$ is a critical point of $\overline{g}$ exactly if both $p$ is a critical point of $g$ modulo $\Phi$ and $\Psi_{p}([p])$ is a critical point of $\overline{g}\circ\Psi_{p}^{-1}$; in this case, the index of $[p]$ is the index of $\Psi_{p}([p])$. Finally, since every point in $\Phi^{-1}(0)$ is tall, M//T{M/\!/T} is oriented by [KT1, Proposition 6.1]. ∎
The aim of Sections 4 and 6 is to show that the hypotheses of Proposition 3.4 hold for non-degenerate singular points.
4. Purely elliptic type
Our goal in this section is to prove that tall critical points of $g$ modulo $\Phi$ that have purely elliptic type satisfy the hypotheses of Proposition 3.4 (see Proposition 4.5 below). An important ingredient is the local classification of such points (see Lemma 4.4).
We begin by giving an important example and describing the symplectic form on (complexity one) local models.
Example 4.1.
Fix $\ell\in\{0,1,\dots,n\}.$ Let $\omega_{0}\in\Omega^{2}(T^{*}(S^{1})^{n-\ell}\times{\mathbb{C}}^{\ell})$ be the sum of the pullbacks of the canonical symplectic form on $T^{*}(S^{1})^{n-\ell}$ and the standard symplectic form on $\mathbb{C}$$^{\ell}$. Let $\Phi_{0}\colon T^{*}(S^{1})^{n-\ell}\times{\mathbb{C}}^{\ell}\simeq(S^{1})^{n-\ell}\times\left({\mathbb{R}}^{n-\ell}\right)^{*}\times{\mathbb{C}}^{\ell}\to({\mathbb{R}}^{n})^{*}$ be the moment map for the standard $(S^{1})^{n}=(S^{1})^{n-\ell}\times(S^{1})^{\ell}$-action on $T^{*}(S^{1})^{n-\ell}\times{\mathbb{C}}^{\ell}$ given by $\Phi_{0}(s,a,w)=(a,\frac{1}{2}|w|^{2})$, where here $\frac{1}{2}|w|^{2}:=\frac{1}{2}(|w_{1}|^{2},\ldots,|w_{\ell}|^{2})$. Then $(T^{*}(S^{1})^{n-\ell}\times{\mathbb{C}}^{\ell},\omega_{0},{\mathbb{R}}^{n},\Phi_{0})$ is an integrable system; moreover, $p=(s,a,w)$ is singular exactly if $\prod_{i}w_{i}=0$, in which case $p$ has purely elliptic type.
Remark 4.2.
Let $Y=T\times_{H}{\mathfrak{h}}^{\circ}\times{\mathbb{C}}^{h+1}$ be a complexity one local model, where $H$ acts on ${\mathbb{C}}^{h+1}$ via $\rho:H\to(S^{1})^{h+1}$ with homogeneous moment map $\Phi_{H}\colon{\mathbb{C}}^{h+1}\to{\mathfrak{h}}^{*}$. Let $i^{*}\colon{\mathfrak{t}}^{*}\to{\mathfrak{h}}^{*}$ be the projection induced by the inclusion $H\hookrightarrow T$. Equip $T^{*}T\times{\mathbb{C}}^{h+1}$ with the symplectic form given by the sum of the pullbacks of the canonical symplectic form on $T^{*}T$ and the standard symplectic form on ${\mathbb{C}}^{h+1}$. The $H$-action on $T^{*}T\times{\mathbb{C}}^{h+1}\cong T\times{\mathfrak{t}}^{*}\times{\mathbb{C}}^{h+1}$ given by $h\cdot(t,\beta,w):=(ht,\beta,\rho(t^{-1})w)$ has a moment map $T\times{\mathfrak{t}}^{*}\times{\mathbb{C}}^{h+1}\to{\mathfrak{h}}^{*}$ that takes $(t,\beta,z)$ to $i^{*}(\beta)-\Phi_{H}(z)$. The reduced space at 0 is smooth, and the map $Y\to(T\times{\mathfrak{t}}^{*}\times{\mathbb{C}}^{h+1})/\!/H$ taking $t,\alpha,z]$ to $t,\alpha+\Phi_{H}(z),z]$ is a diffeomorphism. By definition, the symplectic form $\omega_{Y}$ on $Y$ is the pullback of the reduced symplectic form along this diffeomorphism.
Next we define “isomorphism” for the class of integrable systems that we consider.
Definition 4.3.
Let $(M_{i},\omega_{i},{\mathfrak{t}}\times{\mathbb{R}},f_{i}=(\Phi_{i},g_{i}))$ be an integrable system such that $(M_{i},\omega_{i},\Phi_{i})$ is a complexity one $T$-space for $i=1,2$. An isomorphism from $(M_{1},\omega_{1},{\mathfrak{t}}\times{\mathbb{R}},f_{1})$ to $(M_{2},\omega_{2},{\mathfrak{t}}\times{\mathbb{R}},f_{2})$ is a symplectomorphism $\Psi\colon(M_{1},\omega_{1})\to(M_{2},\omega_{2})$ and a diffeomorphism [^2] Let $X,Z$ be manifolds and let $A\subseteq X$. A map $\psi\colon A\to Z$ is smooth if, for every point $a\in A$, there exists an open set $U_{a}\subseteq X$ and a smooth map $\psi_{a}\colon U_{a}\to Z$ extending $\psi$. $\psi\colon f_{1}(M_{1})\subseteq{\mathfrak{t}}^{*}\times{\mathbb{R}}\to f_{2}(M_{2})\subseteq{\mathfrak{t}}^{*}\times{\mathbb{R}}$ of the form
$\psi(\beta,c)=(\tau(\beta),\psi_{{\mathbb{R}}}(\beta,c))\text{ for }(\beta,c)\in f_{1}(M_{1}),$
where $\tau\colon{\mathfrak{t}}^{*}\to{\mathfrak{t}}^{*}$ is a translation, such that $f_{2}\circ\Psi=\psi\circ f_{1}$.
We can now give the local classification for critical points of $g$ modulo $\Phi$ that have purely elliptic type.
Lemma 4.4.
Let $(M,\omega,{\mathfrak{t}}\times{\mathbb{R}},f:=(\Phi,g))$ be an integrable system such that $(M,\omega,\Phi)$ is a complexity one $T$-space. Let $p\in M$ be a critical point of $g$ modulo $\Phi$ that has purely elliptic type. Let $Y=T\times_{H}{\mathfrak{h}}^{\circ}\times{\mathbb{C}}^{h+1}$ be the local model for $p$, where $H$ acts on ${\mathbb{C}}^{h+1}$ via $\rho\colon H\to(S^{1})^{h+1}$. Given $b\in{\mathbb{R}}^{h+1}\smallsetminus\rho_{*}({\mathfrak{h}})$, a neighborhood of $T\cdot p$ in $(M,\omega,{\mathfrak{t}}\times{\mathbb{R}},f=(\Phi,g))$ is isomorphic to a neighborhood of $T\cdot[1,0,0]$ in $(Y,\omega_{Y},{\mathfrak{t}}\times{\mathbb{R}},f_{b}=(\Phi_{Y},g_{b}))$, where $g_{b}\colon Y\to{\mathbb{R}}$ is given by $g_{b}([t,\alpha,z])=\sum_{i=0}^{h}b_{i}|z_{i}|^{2}$.
Proof.
Since $p$ is a critical point of $g$ modulo $\Phi$, ${\operatorname{ker}}D_{p}f={\operatorname{ker}}D_{p}\Phi$, and so the $\left(T\times{\mathbb{R}}\right)$-orbit of $p$ is equal to the $T$-orbit of $p$. In particular, the rank of $D_{p}f$ is nh1n-h-1 and the $\left(T\times{\mathbb{R}}\right)$-orbit of $p$ is compact, where $h$ is the dimension of the stabilizer of $p$ for the $T$-action. By the local normal form for compact orbits of points that have purely elliptic type [E], we may assume that $M$ and the symplectic form are as in Example 4.1, where $\ell=h+1$, that $p=(1,0,0)$, and that there exists a diffeomorphism $\varphi\colon\left({\mathbb{R}}^{n}\right)^{*}\to{\mathfrak{t}}^{*}\times{\mathbb{R}}$ such that $f=\varphi\circ\Phi_{0}$, where $\Phi_{0}$ is also as in Example 4.1.
Since $\varphi\circ\Phi_{0}=f=(\Phi,g)$, there exist an injective Lie group homomorphism $\Lambda\colon T\to(S^{1})^{n}=(S^{1})^{n-h-1}\times(S^{1})^{h+1}$, $\beta\in{\mathfrak{t}}^{*}$, and a smooth map $\varphi_{\mathbb{R}}\colon({\mathbb{R}}^{n})^{*}\to{\mathbb{R}}$ such that $\varphi=(\Lambda^{*}+\beta,\varphi_{{\mathbb{R}}})$, where $\Lambda^{*}$ is the dual of the Lie algebra homomorphism $\Lambda_{*}:{\mathfrak{t}}\to{\mathbb{R}}^{n}$. By replacing $f$ by its pre-composition with the translation by $\beta$, we may assume that $\beta=0$. Hence, the composite $D_{0}\varphi\circ\varphi^{-1}\colon{\mathfrak{t}}^{*}\times{\mathbb{R}}\to{\mathfrak{t}}^{*}\times{\mathbb{R}}$ is the identity on $\mathfrak{t}$$^{*}$. Therefore, by replacing $f$ by the composition $D_{0}\varphi\circ\varphi^{-1}\circ f=D_{0}\varphi\circ\Phi_{0}$, we may assume that $\varphi_{{\mathbb{R}}}$ is linear.
Let $\Lambda_{1}$ and $\Lambda_{2}$ denote the projections of $\Lambda$ onto $(S^{1})^{n-h-1}$ and $(S^{1})^{h+1}$, and let $\Lambda_{1}^{*}$ and $\Lambda^{*}_{2}$ be the dual of the Lie algebra homomorphisms induced by $\Lambda_{1}$ and $\Lambda_{2}$ respectively. Then ${\operatorname{ker}}\,\Lambda_{1}=H$ and $\Lambda_{2}|_{H}=\rho$. Hence, if $b\in{\mathbb{R}}^{h+1}\smallsetminus\rho_{*}({\mathfrak{h}})$, then $(0,b)\in{\mathbb{R}}^{n}\smallsetminus\Lambda_{*}({\mathfrak{t}})$. Thus there exists a linear isomorphism $L$ of $\mathfrak{t}$$^{*}\times{\mathbb{R}}$ that restricts to the identity on $\mathfrak{t}$$^{*}$ so that, by further replacing $f$ by the composition $L\circ f$, we may assume that $\varphi_{{\mathbb{R}}}=(0,2b)$. In particular, $g(s,a,w)=\sum_{i=0}^{h}b_{i}|w_{i}|^{2}$.
By construction, the map given by
$t,\alpha,z]\mapsto[t,\alpha+\Phi_{H}(z),z]$
is a symplectomorphism between the local model $Y$ and the reduced space $(T\times{\mathfrak{t}}^{*}\times{\mathbb{C}}^{h+1})/\!/H$ discussed in Remark 4.2.
By a straightforward computation, the map $T\times{\mathfrak{t}}^{*}\times{\mathbb{C}}^{h+1}\to T\times{\mathfrak{t}}^{*}\times{\mathbb{C}}^{h+1}$ given by $(t,\beta,z)\mapsto\left(t,\beta-\Lambda^{*}_{2}\left(\frac{1}{2}|z_{0}|^{2},\ldots,\frac{1}{2}|z_{h}|^{2}\right),\Lambda_{2}(t)z\right)$ is a symplectomorphism. It induces a symplectomorphism between the reduced space $(T\times{\mathfrak{t}}^{*}\times{\mathbb{C}}^{h+1})/\!/H$ and $T/H\times{\mathfrak{h}}^{\circ}\times{\mathbb{C}}^{h+1}$ that is given by
$[t,\beta,z]\mapsto(tH,\beta-\Lambda^{*}_{2}\left(\frac{1}{2}|z_{0}|^{2},\ldots,\frac{1}{2}|z_{h}|^{2}\right),\Lambda_{2}(t)z).$
Here the symplectic form on $T/H\times{\mathfrak{h}}^{\circ}\times{\mathbb{C}}^{h+1}$ is the sum of the pullbacks of the canonical symplectic form on $T^{*}(T/H)\cong T/H\times{\mathfrak{h}}^{\circ}$ and of the standard symplectic form on ${\mathbb{C}}^{h+1}$.
By a dimension count, $\Lambda_{1}$ induces an isomorphism between $T/H$ and $(S^{1})^{n-h-1}$. Hence it induces a symplectomorphism between $T/H\times{\mathfrak{h}}^{\circ}\times{\mathbb{C}}^{h+1}$ and $(S^{1})^{n-h-1}\times\left({\mathbb{R}}^{n-h-1}\right)^{*}\times{\mathbb{C}}^{h+1}$ that is given by
$(tH,\alpha,w)\mapsto(\Lambda_{1}(t),(\Lambda^{*}_{1})^{-1}(\alpha),w).$
Composing these maps, we obtain a symplectomorphism $\Psi\colon Y\to(S^{1})^{n-h-1}\times\left({\mathbb{R}}^{n-h-1}\right)^{*}\times{\mathbb{C}}^{h+1}$. Since $\Lambda_{2}|_{H}=\rho$, it is given by
$\Psi([t,\alpha,z])=\left(\Lambda_{1}(t),(\Lambda_{1}^{*})^{-1}\left(\alpha-\mathrm{pr}_{{\mathfrak{h}}^{\circ}}\left(\Lambda^{*}_{2}\left(\frac{1}{2}|z_{0}|^{2},\ldots,\frac{1}{2}|z_{h}|^{2}\right)\right)\right),\Lambda_{2}(t)z\right).$
Here $\mathrm{pr}_{{\mathfrak{h}}^{\circ}}\colon{\mathfrak{t}}^{*}\to{\mathfrak{h}}^{\circ}$ denotes the projection induced by the fixed choice of metric on $\mathfrak{t}$}. Clearly this map is $T$-equivariant and so $\Phi\circ\Psi$ is a moment map for the $T$-action on $Y$. Since $\Phi\circ\Psi(1,0,0)=0$, it follows that $\Phi\circ\Psi$ is the homogeneous moment map $\Phi_{Y}$. Moreover, it is easy check that $g\circ\Psi=g_{b}$. By construction, the integrable system $(Y,\omega_{Y},{\mathfrak{t}}\times{\mathbb{R}},f\circ\Psi=(\Phi_{Y},g_{b}))$ is isomorphic to the original one, as desired. ∎
We are finally ready to prove the main result of this section.
Proposition 4.5.
Let $(M,\omega,{\mathfrak{t}}\times{\mathbb{R}},f:=(\Phi,g))$ be an integrable system such that $(M,\omega,\Phi)$ is a complexity one $T$-space. Let $p\in\Phi^{-1}(0)\cap g^{-1}(0)$ be a tall critical point of $g$ modulo $\Phi$ that has purely elliptic type. After possibly replacing $g$ by g-g, there exist an open neighborhood $U$ of $[p]$ in M//T{M/\!/T} and a map $\Psi\colon U\to{\mathbb{R}}^{2}$ taking $[p]$ to $(0,0)$ satisfying:
  1. (1)
    $\Psi\colon U\to\Psi(U)$ is a homeomorphism onto an open set;
  2. (2)
    $\Psi$ restricts to a diffeomorphism on $U\smallsetminus\{[p]\}$; and
  3. (3)
    $(\overline{g}\circ\Psi^{-1})(x,y)=x^{2}+y^{2}$ for all $(x,y)\in\Psi(U)$.
Proof.
Since $p$ has tall local model $Y=T\times_{H}{\mathfrak{h}}^{\circ}\times{\mathbb{C}}^{h+1}$, Remark 2.3 implies that $(1,\ldots,1)\notin\rho_{*}({\mathfrak{h}})$, where $\rho\colon H\to(S^{1})^{h+1}$ is the symplectic slice representation at $p$. Hence, by Lemma 4.4 and the Morse Lemma, after possibly replacing $g$ by g-g, we may assume that $M$ is a local model $Y=T\times_{H}{\mathfrak{h}}^{\circ}\times{\mathbb{C}}^{h+1}$, that $\Phi=\Phi_{Y}$, that $g([t,\alpha,z])=|z|^{2}$, and that $p=[1,0,0]$. Let $N$ be the degree of the defining polynomial $P\colon{\mathbb{C}}^{h+1}\to{\mathbb{C}}$. By Lemma 2.7, the map $\overline{P}\colon{Y/\!/T}\to{\mathbb{C}}$ is a homeomorphism; moreover, it restricts to a diffeomorphism from $({Y/\!/T})\smallsetminus\{[t,0,0]\}$ to ${\mathbb{C}}\smallsetminus\{0\}$. By Lemma 2.9, there exists a positive real number $C$ such that $g([t,\alpha,z])=C|P(z)|^{\frac{2}{N}}$ for all $t,\alpha,z]\in\Phi_{Y}^{-1}(0)$. Let $S\colon{\mathbb{C}}\to{\mathbb{C}}$ be the homeomorphism that, in polar coordinates, takes $(r,\theta)$ to $(\sqrt{C}\,r^{\frac{1}{N}},\theta)$; it restricts to a diffeomorphism ${\mathbb{C}}\smallsetminus\{0\}\to{\mathbb{C}}\smallsetminus\{0\}$. Define $\Psi\colon{Y/\!/T}\to{\mathbb{C}}\simeq{\mathbb{R}}^{2}$ by $\Psi([t,\alpha,z])=S(P(z))$. Since
$|\Psi([t,\alpha,z])|^{2}=C|P(z)|^{\frac{2}{N}}=\overline{g}([t,\alpha,z]),$
for all $t,\alpha,z]\in{Y/\!/T}$, the map $\Psi$ satisfies properties (1) – (3). ∎
Remark 4.6.
We observe that, by Lemma 2.8, if $N=1$ the map $\Psi$ constructed in the above proof is a diffeomorphism.
5. The semitoric case
The goal of this section is to prove that focus-focus points satisfy the hypotheses of Proposition 3.4 (see Lemma 5.4 below). This allows us to provide a short and direct new proof of Vu Ngoc’s theorem that the fibers of semitoric systems are connected [VN] – see Theorem 5.7. In both proofs, the key arguments follow from Morse theory; the most significant difference is that we construct a new smooth structure on the reduced spaces, while he uses the natural smooth structure on $M$.
Definition 5.1.
An integrable system $\left(M,\omega,{\mathfrak{t}}\times{\mathbb{R}},f=\left(\Phi,g\right)\right)$ is semitoric if
  • $\left(M,\omega,\Phi\right)$ is a 4-dimensional complexity one $T$-space, and
  • every singular point of $\left(M,\omega,{\mathfrak{t}}\times{\mathbb{R}},f\right)$ is non-degenerate with no hyperbolic blocks.
Note that, by definition, every singular point of a semitoric system either has purely elliptic type or is focus-focus, i.e., it has one focus-focus block. We begin with an important example of a semitoric system containing one focus-focus point.
Example 5.2.
Let $Y={\mathbb{C}}^{2}$, let $\omega_{Y}$ be the standard symplectic form, and set $\Phi_{Y}(z_{1},z_{2}):=\frac{1}{2}\left(|z_{1}|^{2}-|z_{2}|^{2}\right)$ and $g_{Y}(z_{1},z_{2}):={\mathfrak{I}}(z_{1}z_{2})={\mathfrak{I}}(P(z_{1},z_{2}))$, where $P\colon Y\to{\mathbb{C}}$ is the defining polynomial for the $S^{1}$-action induced by $\Phi_{Y}$ (cf. Example 2.6).
Locally, near a focus-focus point, every semitoric system is modeled on Example 5.2.
Proposition 5.3.
Let $(M,\omega,{\mathfrak{t}}\times{\mathbb{R}},f=(\Phi,g))$ be an integrable system such that $(M,\omega,\Phi)$ is a 4-dimensional complexity one $T$-space. Assume that $p\in M$ is focus-focus. After identifying $T$ with $S^{1}$, a neighborhood of $p$ is isomorphic to a neighborhood of the origin in the semitoric system $(Y,\omega_{Y},{\mathbb{R}}\times{\mathbb{R}},f_{Y})$ described above.
Proof.
Identify $T$ with $S^{1}$. By the local normal form for focus-focus points, we may assume that the integrable system is $(Y,\omega_{Y},{\mathbb{R}}\times{\mathbb{R}},\varphi\circ f_{Y})$ and $p$ is the origin, where $(Y,\omega_{Y},{\mathbb{R}}\times{\mathbb{R}},f_{Y}=(\Phi_{Y},g_{Y}))$ is as in Example 5.2 and $\varphi\colon{\mathbb{R}}\times{\mathbb{R}}\to{\mathbb{R}}\times{\mathbb{R}}$ is a diffeomorphism [C, VW]. Since by assumption the first component of $\varphi\circ f_{Y}$ induces an effective $S^{1}$-action on $Y$, and since the fibers of $f_{Y}$ are not compact, up to a translation $\Phi_{Y}$ and the first component of $\varphi\circ f_{Y}$ are equal up to sign (cf. [HSS, Proposition 3.9]). Since the integrable systems $(Y,\omega_{Y},{\mathbb{R}}\times{\mathbb{R}},f_{Y}=(\Phi_{Y},g_{Y}))$ and $(Y,\omega_{Y},{\mathbb{R}}\times{\mathbb{R}},(-\Phi_{Y},g_{Y}))$ are isomorphic, the claim follows. ∎
Hence, focus-focus points satisfy the hypotheses of Proposition 3.4.
Lemma 5.4.
Let $(M,\omega,{\mathfrak{t}}\times{\mathbb{R}},f=(\Phi,g))$ be an integrable system such that $(M,\omega,\Phi)$ is a 4-dimensional complexity one $T$-space. Assume that $p\in\Phi^{-1}(0)\cap g^{-1}(0)$ is focus-focus. Then there exist an open neighborhood $U$ of $p]\in{M/\!/T}$ and a map $\Psi\colon U\to{\mathbb{R}}^{2}$ taking $[p]$ to $(0,0)$ satisfying:
  1. (1)
    $\Psi\colon U\to\Psi(U)$ is a homeomorphism onto an open set;
  2. (2)
    $\Psi$ restricts to a diffeomorphism on $U\smallsetminus\{[p]\}$; and
  3. (3)
    $(\overline{g}\circ\Psi^{-1})(x,y)=y$ for all $(x,y)\in\Psi(U)$.
Proof.
By Proposition 5.3 we may assume that the integrable system is the system $(Y,\omega_{Y},{\mathbb{R}}\times{\mathbb{R}},f_{Y}:=(\Phi_{Y},g_{Y}))$ described in Example 5.2, and that $p$ is the origin. Set $\Psi[z_{1},z_{2}]:=z_{1}z_{2}=P(z_{1},z_{2})$. Then the claim follows immediately from Lemma 2.7. ∎
We are now almost ready to give our proof of Vu Ngoc’s theorem on the connectedness of the fibers of semitoric systems; however, we also need the following important result, which is proved in [At, GS2, LMTW].
Theorem 5.5.
Let $T$ act on a connected symplectic manifold $(M,\omega)$ with moment map $\Phi\colon M\to{\mathfrak{t}}^{*}$. If $\Phi$ is proper, then every fiber of $\Phi$ is connected and $\Phi$ is open as a map to $\Phi(M)$.
We also need the following proposition. Atiyah proved the compact case (see [At]), and used it to prove the theorem above; it was later extended to the proper case; see [PRV, Proposition 4.5]. (For completeness, we provide a variation on the arguments in [PRV] in Section 7.)
Proposition 5.6.
Let $N$ be a connected manifold and let $\overline{g}\colon N\to{\mathbb{R}}$ be a proper Morse-Bott function. If no critical point of $\overline{g}$ has index or coindex one, then the fiber $\overline{g}^{-1}(c)$ is connected for all $c\in{\mathbb{R}}$.
We observe that the result below is a special case of Theorem 1.6, and the proof below illustrates our general strategy.
Theorem 5.7 (Vũ Ngoc).
Let $(M,\omega,{\mathfrak{t}}\times{\mathbb{R}},f=(\Phi,g))$ be a semitoric system such that $\Phi$ is proper. Then the fiber $f^{-1}(\beta,c)=\Phi^{-1}(\beta)\cap g^{-1}(c)$ is connected for all $(\beta,c)\in{\mathfrak{t}}^{*}\times{\mathbb{R}}$.
Proof.
Fix $\beta\in{\mathfrak{t}}^{*}$. Assume without loss of generality that $\beta=0$. Since $\Phi$ is proper, the reduced space M//T{M/\!/T} is compact and connected by Theorem 5.5.
If the reduced space M//T{M/\!/T} is a single point then obviously $\Phi^{-1}(0)$ consists of a single $T$-orbit. Hence, the fiber $f^{-1}(0,c)$ is connected for all $c\in{\mathbb{R}}$.
Otherwise, since the orbit of a short point in $\Phi^{-1}(0)$ is an isolated point in $\Phi^{-1}(0)/T$, every point in $\Phi^{-1}(0)$ is tall. If $p\in\Phi^{-1}(0)$ is a critical point of $g$ modulo $\Phi$, then $p$ is also a singular point of $\left(M,\omega,{\mathfrak{t}}\times{\mathbb{R}},f\right)$. So, by assumption, $p$ either has purely elliptic type or is focus-focus. By Propositions 3.4 and 4.5, and Lemma 5.4, M//T{M/\!/T} can be given the structure of a smooth closed, connected, oriented surface such that $\overline{g}:{M/\!/T}\to{\mathbb{R}}$ is a Morse function with no critical point of index 1. By Proposition 5.6, the fiber $\overline{g}^{-1}(c)$ is connected for all $c\in{\mathbb{R}}$. Since $T$ is connected, it follows that $f^{-1}(0,c)$ is also connected. ∎
6. Non-degenerate singular points
Let $\left(M,\omega,{\mathfrak{t}}\times{\mathbb{R}},f:=\left(\Phi,g\right)\right)$ be an integrable system such that $\left(M,\omega,\Phi\right)$ is a complexity one $T$-space. Let $p\in\Phi^{-1}(0)$ be a tall non-degenerate singular point. By Lemma 3.2 and Proposition 4.5, if $p$ has purely elliptic type, then we can put a smooth structure on M//T{M/\!/T} near $[p]$ so that $\overline{g}$ is locally Morse, and hence $p$ satisfies the hypotheses of Proposition 3.4. The aim of this section is to extend this result to the remaining cases (see Proposition 6.3). In particular, if $p$ either has a focus-focus block, or has a hyperbolic block and disconnected $T$-stabilizer, then $[p]$ is a regular point of $\overline{g}$, while if $p$ has a hyperbolic block and connected $T$-stabilizer, then $[p]$ is a singular point with index one. We start by recalling some facts about non-degenerate singular points.
Remark 6.1.
Let $p$ be a point in an integrable system $(M,\omega,V,f)$ and let $W_{0}\subseteq V$ be the Lie algebra of the stabilizer of $p$. Let $\mathcal{O}$ be the $V$-orbit through $p$. The linearized action at $p$ induces a Lie algebra homomorphism $W_{0}\to\mathfrak{sp}((T_{p}\mathcal{O})^{\omega}/T_{p}\mathcal{O})$. Given $\xi\in W_{0}$, we can calculate the associated symplectic endomorphism as follows: Let $f_{\mathrm{lin}}$ be the homogeneous moment map for the linearized action. We view $\langle f_{\mathrm{lin}},\xi\rangle$ as a homogeneous quadratic polynomial on $(T_{p}\mathcal{O})^{\omega}/T_{p}\mathcal{O}$. We recall that there is a standard isomorphism between homogeneous quadratic polynomials and $\mathfrak{sp}((T_{p}\mathcal{O})^{\omega}/T_{p}\mathcal{O})$ that, upon choosing a standard basis for the symplectic vector space $(T_{p}\mathcal{O})^{\omega}/T_{p}\mathcal{O}$, sends a symmetric matrix $A$ to $JA$, where $J$ is the standard complex structure on $(T_{p}\mathcal{O})^{\omega}/T_{p}\mathcal{O}$. The above Lie algebra homomorphism sends $\xi$ to $J\langle f_{\mathrm{lin}},\xi\rangle$.
If, in addition, $p$ is non-degenerate, then for a generic element $\xi$ in the Lie algebra of $W$, the eigenvalues of the above endomorphism of $(T_{p}\mathcal{O})^{\omega}/T_{p}\mathcal{O}$ can be characterized as follows: If $f_{\mathrm{lin}}$ can be written as the product of $k_{e}$ elliptic blocks, $k_{h}$ hyperbolic blocks and $k_{f}$ focus-focus blocks, then the above endomorphism has exactly $2k_{e}$ distinct non-zero imaginary eigenvalues, $2k_{h}$ distinct non-zero real eigenvalues, and $4k_{f}$ distinct non-zero complex eigenvalues that are neither real nor imaginary. In particular, the numbers $k_{e},k_{h}$ and $k_{f}$ are a well-defined invariant of $p$.
Before proceeding to the tall case, we start by analyzing short points.
Lemma 6.2.
Let $\left(M,\omega,{\mathfrak{t}}\times{\mathbb{R}},f:=\left(\Phi,g\right)\right)$ be an integrable system such that $\left(M,\omega,\Phi\right)$ is a complexity one $T$-space. If $p\in M$ is short and non-degenerate, then $p$ has purely elliptic type.
Proof.
For simplicity, assume that $g(p)=0$. By the Marle-Guillemin-Sternberg local normal form theorem, we may assume that $M$ is a local model $Y=T\times_{H}{\mathfrak{h}}^{\circ}\times{\mathbb{C}}^{h+1}$, that $\Phi=\Phi_{Y}$, and that $p=[1,0,0]$. Let $j\colon{\mathbb{C}}^{h+1}\to T\times_{H}{\mathfrak{h}}^{\circ}\times{\mathbb{C}}^{h+1}$ denote the inclusion $z\mapsto[1,0,z]$. By Lemma 3.2, $p$ is a critical point of $g$ modulo $\Phi$. Hence, the $(T\times{\mathbb{R}})$-orbit $\mathcal{O}$ through $p$ coincides with the $T$-orbit $\mathfrak{O}$ through $p$ and the Taylor polynomial $T^{1}_{0}(j^{*}g)$ of degree one at $0\in{\mathbb{C}}^{h+1}$ vanishes. Thus we may identify $(T_{p}\mathcal{O})^{\omega}/T_{p}\mathcal{O}=(T_{p}\mathfrak{O})^{\omega}/T_{p}\mathfrak{O}$ with ${\mathbb{C}}^{h+1}$ so that the homogeneous moment map for the linearized action at $p$ is the Taylor polynomial of degree two $T^{2}_{0}(j^{*}f)=(T^{2}_{0}(j^{*}\Phi),T^{2}_{0}(j^{*}g))=(\Phi_{H},T^{2}_{0}(j^{*}g))$, where $\Phi_{H}\colon{\mathbb{C}}^{h+1}\to{\mathfrak{h}}^{*}$ is the homogeneous moment map for the $H$-action.
By [KT1, Lemma 5.4], $\Phi$ is proper; hence, so is $(\Phi_{H},T_{0}^{2}(j^{*}g))$. If $p$ has a hyperbolic or a focus-focus block, then the homogeneous moment map for the linearized $(H\times{\mathbb{R}})$-action on ${\mathbb{C}}^{h+1}$ is not proper. Hence, $p$ has purely elliptic type. ∎
We can now state the main result of this section.
Proposition 6.3.
Let $\left(M,\omega,{\mathfrak{t}}\times{\mathbb{R}},f:=\left(\Phi,g\right)\right)$ be an integrable system such that $\left(M,\omega,\Phi\right)$ is a complexity one $T$-space. Assume that $p\in\Phi^{-1}(0)\cap g^{-1}(0)$ is non-degenerate and tall; let $N$ be the degree of the defining polynomial of the local model for $p$. Then $p$ has at most one non-elliptic block; furthermore,
  1. (1)
    If $p$ has a hyperbolic block and connected $T$-stabilizer, then $N=1$. Moreover, there exist an open neighborhood $U$ of $p]\in{M/\!/T}$ and a map $\Psi\colon U\to{\mathbb{R}}^{2}$ taking $[p]$ to $(0,0)$ satisfying:
    1. (a)
      $\Psi\colon U\to\Psi(U)$ is a diffeomorphism onto an open set; and
    2. (b)
      $(\overline{g}\circ\Psi^{-1})(x,y)=x^{2}-y^{2}$ for all $(x,y)\in\Psi(U)$.
  2. (2)
    If $p$ either has a focus-focus block, or has a hyperbolic block and disconnected $T$-stabilizer, then $N=2$. Moreover, there exist an open neighborhood $U$ of $p]\in{M/\!/T}$ and a map $\Psi\colon U\to{\mathbb{R}}^{2}$ taking $[p]$ to $(0,0)$ satisfying:
    1. (a)
      $\Psi\colon U\to\Psi(U)$ is a homeomorphism onto an open set;
    2. (b)
      $\Psi$ restricts to a diffeomorphism on $U\smallsetminus\{[p]\}$; and
    3. (c)
      $(\overline{g}\circ\Psi^{-1})(x,y)=y$ for all $(x,y)\in\Psi(U)$.
Proof.
By the Marle-Guillemin-Sternberg local normal form theorem, we may assume that $M$ is a tall local model $Y=T\times_{H}{\mathfrak{h}}^{\circ}\times{\mathbb{C}}^{h+1}$, that $\Phi=\Phi_{Y}$, and that $p=[1,0,0]$. Let $j\colon{\mathbb{C}}^{h+1}\to T\times_{H}{\mathfrak{h}}^{\circ}\times{\mathbb{C}}^{h+1}$ denote the inclusion $z\mapsto[1,0,z]$.
By Lemma 3.2, every regular point of $g$ modulo $\Phi$ has purely elliptic type. Therefore, we may assume that $p$ is a critical point of $g$ modulo $\Phi$. As in the proof of Lemma 6.2, this implies that the Taylor polynomial of degree one $T^{1}_{0}(j^{*}g)$ is identically zero, and we may identify $(T_{p}\mathcal{O})^{\omega}/T_{p}\mathcal{O}$ with ${\mathbb{C}}^{h+1}$ so that the homogeneous moment map for the linearized action at $p$ is the Taylor polynomial of degree two $(\Phi_{H},T^{2}_{0}(j^{*}g))$.
Let $P\colon{\mathbb{C}}^{h+1}\to{\mathbb{C}}$ be the defining polynomial of $Y$. Depending on the degree $N$ of $P$, we may assume that we are in one of the following four cases (cf. Examples 2.5 and 2.6).
Case 1: $\mathbf{N=1}$ and $\mathbf{P(z)=z_{0}}$. In this case, ${\mathbb{C}}^{h+1}\simeq{\mathbb{C}}\times{\mathbb{C}}^{h}$ and $H\simeq(S^{1})^{h}$, where $(S^{1})^{h}$ acts on ${\mathbb{C}}^{h}$ in the standard fashion and trivially on $\mathbb{C}$}. Thus the span of $\{\Phi_{H}^{\xi}\mid\xi\in{\mathfrak{h}}\}$ equals the span of $\{|z_{k}|^{2}\colon k=1,\ldots,h\}.$ Moreover, since the Taylor polynomial $T^{2}_{0}(j^{*}g)$ is $H$-invariant, by Remark 2.12 it lies in the span of $\{{\mathfrak{R}}(z_{0}^{2}),{\mathfrak{I}}(z_{0}^{2}),|z_{k}|^{2}\colon k=0,\ldots,h\}$. Therefore there exist constants $a_{x},a_{y},a_{\rho}\in{\mathbb{R}}$ and $\xi\in{\mathfrak{h}}$ such that
(6.4) $T^{2}_{0}(j^{*}g)(z)=a_{x}{\mathfrak{R}}(z_{0}^{2})+a_{y}{\mathfrak{I}}(z_{0}^{2})+a_{\rho}|z_{0}|^{2}+\Phi_{H}^{\xi}(z)\ \forall z\in{\mathbb{C}}^{h+1}.$
Moreover, a generic element in the span of $\{T^{2}_{0}(j^{*}g),\Phi_{H}^{\xi}\mid\xi\in{\mathfrak{h}}\}$ is a multiple of
$a_{x}{\mathfrak{R}}(z^{2}_{0})+a_{y}{\mathfrak{I}}(z^{2}_{0})+a_{\rho}|z_{0}|^{2}+\sum\limits_{k=1}^{h}c_{k}|z_{k}|^{2},$
where $c_{k}\in{\mathbb{R}}\smallsetminus\{0\}$ for $k=1,\ldots,h$. The eigenvalues of the corresponding symplectic endomorphism are given by
$\pm 2\sqrt{a_{x}^{2}+a_{y}^{2}-a_{\rho}^{2}},\,\,\pm 2c_{1}\sqrt{-1},\,\,\ldots,\,\,\pm 2c_{h}\sqrt{-1}.$
Hence, since $p$ is non-degenerate, Remark 6.1 implies that $a_{x}^{2}+a_{y}^{2}-a_{\rho}^{2}\neq 0$. Moreover, if $a_{x}^{2}+a_{y}^{2}-a_{\rho}^{2}>0$, then $p$ has one hyperbolic block and the remaining blocks are elliptic; otherwise, $p$ has purely elliptic type.
By Lemma 2.8, the inclusion $i\colon{\mathbb{C}}\to Y$ sending $z_{0}$ to $1,0,(z_{0},0,\ldots,0)]$ induces a diffeomorphism $\overline{i}\colon{\mathbb{C}}\to{Y/\!/T}$. By definition of the inclusion $j$ and by (6.4),
$T^{2}_{0}(\overline{i}^{*}\overline{g})(z_{0})=a_{x}{\mathfrak{R}}(z_{0}^{2})+a_{y}{\mathfrak{I}}(z_{0}^{2})+a_{\rho}|z_{0}|^{2}\quad\forall z_{0}\in{\mathbb{C}}.$
Hence, the zero set of the map ${Y/\!/T}\to{\mathbb{R}}$ taking $t,\alpha,z]\in{Y/\!/T}$ to $T^{2}_{0}(j^{*}g)(z)$ is homeomorphic to the set $\{(x,y)\in{\mathbb{R}}^{2}\mid xy=0\}$ if $p$ has a hyperbolic block, and is a single point if $p$ has purely elliptic type.
Finally, assume that $p$ has a hyperbolic block. Then 0 is a non-degenerate critical point of $\overline{i}^{*}\overline{g}$ of index 1. Hence, by the Morse lemma, there exists an open neighborhood $V$ of $(0,0)\in{\mathbb{R}}^{2}$ and a map $\psi\colon V\to{\mathbb{C}}$ taking $(0,0)$ to 0 that is a diffeomorphism onto an open set such that $(\overline{i}\circ\psi)^{*}\overline{g}(x,y)=x^{2}-y^{2}$ for all $(x,y)\in V$. Setting $U:=(\overline{i}\circ\psi)(V)$ and $\Psi\colon U\to{\mathbb{R}}^{2}$ to be $(\overline{i}\circ\psi)^{-1},$ the result follows.
Case 2: $\mathbf{N=2}$ and $\mathbf{P(z)=z_{0}^{2}}$. In this case, ${\mathbb{C}}^{h+1}\simeq{\mathbb{C}}\times{\mathbb{C}}^{h}$ and $H\simeq\mathbb{Z}_{2}\times(S^{1})^{h}$, where $\mathbb{Z}_{2}$ acts on $\mathbb{C}$} by multiplication with 1-1 and trivially on ${\mathbb{C}}^{h}$, and $(S^{1})^{h}$ acts on ${\mathbb{C}}^{h}$ in the standard fashion and trivially on $\mathbb{C}$}. Thus the span of $\{\Phi_{H}^{\xi}\mid\xi\in{\mathfrak{h}}\}$ equals the span of $\{|z_{k}|^{2}\colon k=1,\ldots,h\}$. Moreover, since the Taylor polynomial $T^{2}_{0}(j^{*}g)$ is $H$-invariant, by Remark 2.12 it lies in the span of $\{{\mathfrak{R}}(z_{0}^{2}),{\mathfrak{I}}(z_{0}^{2}),|z_{k}|^{2}\colon k=0,\ldots,h\}$. Hence, there exist constants $a_{x},a_{y},a_{\rho}\in{\mathbb{R}}$ and $\xi\in{\mathfrak{h}}$ such that
(6.5) $T^{2}_{0}(j^{*}g)(z)=a_{x}{\mathfrak{R}}(z_{0}^{2})+a_{y}{\mathfrak{I}}(z_{0}^{2})+a_{\rho}|z_{0}|^{2}+\Phi^{\xi}(z)\ \forall z\in{\mathbb{C}}^{h+1}.$
As in the proof of Case 1, since $p$ is non-degenerate, Remark 6.1 implies that $a_{x}^{2}+a_{y}^{2}-a_{\rho}^{2}\neq 0$. Moreover, if $a_{x}^{2}+a_{y}^{2}-a_{\rho}^{2}>0$, then $p$ has one hyperbolic block and the remaining blocks are all elliptic; otherwise, $p$ has purely elliptic type.
By Lemma 2.8, the inclusion $i\colon{\mathbb{C}}\hookrightarrow Y$ sending $z_{0}$ to $1,0,(z_{0},0,\ldots,0)]$ induces a diffeomorphism $\overline{i}$ between ${\mathbb{C}}/\!/{\mathbb{Z}}_{2}={\mathbb{C}}/{\mathbb{Z}}_{2}$ and Y//T{Y/\!/T}. By definition of the inclusion $j$ and by (6.5),
$T^{2}_{0}(i^{*}g)(z_{0})=a_{x}{\mathfrak{R}}(z_{0}^{2})+a_{y}{\mathfrak{I}}(z_{0}^{2})+a_{\rho}|z_{0}|^{2}\quad\forall z_{0}\in{\mathbb{C}}.$
This implies that the zero set of the map ${Y/\!/T}\to{\mathbb{R}}$ taking $t,\alpha,z]\in{Y/\!/T}$ to $T^{2}_{0}(j^{*}g)(z)$ is homeomorphic to $\mathbb{R}$} exactly if $p$ has a hyperbolic block, and is a single point if $p$ has purely elliptic type.
Finally, assume that $p$ has a hyperbolic block. Then 0 is a non-degenerate critical point of $i^{*}g$ of index 1. Hence, by the equivariant Morse lemma, there exists an open neighborhood $V$ of $(0,0)\in{\mathbb{R}}^{2}$ and a map $\psi\colon V\to{\mathbb{C}}$ taking $(0,0)$ to 0 that is an equivariant diffeomorphism onto an open set such that $(i\circ\psi)^{*}g(x,y)=xy$ for all $(x,y)\in V$. Let $\pi\colon{\mathbb{R}}^{2}/{\mathbb{Z}}_{2}\to{\mathbb{R}}^{2}$ be the map that sends $[x,y]$ to $(xy,x^{2}-y^{2})$; it is a homeomorphism onto ${\mathbb{R}}^{2}$ and a diffeomorphism on the complement on $(0,0)$. Setting $U:=(\overline{i}\circ\overline{\psi})(V/{\mathbb{Z}}_{2})$ and $\Psi\colon U\to{\mathbb{R}}^{2}$ to be $\pi\circ(\overline{i}\circ\overline{\psi})^{-1},$ the result follows.
Case 3: $\mathbf{N=2}$ and $\mathbf{P(z)=z_{0}z_{1}}$. In this case, ${\mathbb{C}}^{h+1}\simeq{\mathbb{C}}^{2}\times{\mathbb{C}}^{h-1}$ and $H\simeq S^{1}\times(S^{1})^{h-1}$, where $S^{1}$ acts on ${\mathbb{C}}^{2}$ with weights $\pm 1$ and trivially on ${\mathbb{C}}^{h-1}$, and $(S^{1})^{h-1}$ acts on ${\mathbb{C}}^{h-1}$ in the standard fashion and trivially on ${\mathbb{C}}^{2}$. Hence, the span of $\{\Phi_{H}^{\xi}\mid\xi\in{\mathfrak{h}}\}$ equals the span of $\{|z_{k}|^{2}\colon k=2,\ldots,h\}$ and $\Phi_{1}(z):=\frac{1}{2}(|z_{0}|^{2}-|z_{1}|^{2})$. Moreover, since the Taylor polynomial $T^{2}_{0}(j^{*}g)$ is $H$-invariant, by Remark 2.12 it lies in the span of $\{{\mathfrak{R}}(z_{0}z_{1}),{\mathfrak{I}}(z_{0}z_{1}),|z_{k}|^{2}\colon k=0,\ldots,h\}$. Therefore, there exist constants $a_{x},a_{y},a_{\rho}\in{\mathbb{R}}$ and $\xi\in{\mathfrak{h}}$ such that
(6.6) $T^{2}_{0}(j^{*}g)(z)=a_{x}{\mathfrak{R}}(z_{0}z_{1})+a_{y}{\mathfrak{I}}(z_{0}z_{1})+a_{\rho}(|z_{0}|^{2}+|z_{1}|^{2})+\Phi^{\xi}(z)$
for all $z\in{\mathbb{C}}^{h+1}$. Moreover, a generic element in the span of $\{T^{2}_{0}(j^{*}g),\Phi_{H}^{\xi}\mid\xi\in{\mathfrak{h}}\}$ is a multiple of
$a_{x}{\mathfrak{R}}(z_{0}z_{1})+a_{y}{\mathfrak{I}}(z_{0}z_{1})+a_{\rho}(|z_{0}|^{2}+|z_{1}|^{2})+c_{1}\Phi_{1}+\sum\limits_{k=2}^{h}c_{k}|z_{k}|^{2},$
where $c_{j}\in{\mathbb{R}}\smallsetminus\{0\}$ for $j=1,\ldots,h$. The eigenvalues of the corresponding symplectic endomorphism are given by
$\pm 2c_{1}\sqrt{-1}\pm\sqrt{a_{x}^{2}+a_{y}^{2}-4a_{\rho}^{2}}\,,\,\pm 2c_{2}\sqrt{-1},\,\ldots,\,\pm 2c_{h}\sqrt{-1},$
where the signs in the first term are independent. Hence, since $p$ is non-degenerate, Remark 6.1 implies that $a_{x}^{2}+a_{y}^{2}-4a_{\rho}^{2}\neq 0$. If $a_{x}^{2}+a_{y}^{2}-4a_{\rho}^{2}>0$, then $p$ has one focus-focus block, and the remaining blocks are elliptic; otherwise $p$ has purely elliptic type. It follows from (6.6), and Lemmas 2.7 and 2.9, that the defining polynomial identifies the map ${Y/\!/T}\to{\mathbb{R}}$ taking $t,\alpha,z]\in{Y/\!/T}$ to $T^{2}_{0}(j^{*}g)(z)$ with the map $\mathbb{C}$$\to{\mathbb{R}}$ taking $w\in{\mathbb{C}}$ to $a_{x}{\mathfrak{R}}(w)+a_{y}{\mathfrak{I}}(w)+2a_{\rho}|w|$. Hence, the zero set of the map ${Y/\!/T}\to{\mathbb{R}}$ taking $t,\alpha,z]\in{Y/\!/T}$ to $T^{2}_{0}(j^{*}g)(z)$ is homeomorphic to $\mathbb{R}$} exactly if $p$ has a focus-focus block, and is a single point if $p$ has purely elliptic type.
Finally, assume that $p$ has a focus-focus block. Let $i\colon{\mathbb{C}}^{2}\hookrightarrow Y$ be the inclusion sending $(z_{0},z_{1})$ to $1,0,(z_{0},z_{1},0,\ldots,0)]$. We observe that $i^{*}\omega_{Y}$ is the standard symplectic form on ${\mathbb{C}}^{2}$; see Remark 4.2. Moreover, since $i({\mathbb{C}}^{2})$ is invariant under the $S^{1}$-action generated by $\Phi_{1}$, the functions $i^{*}\Phi_{1}$ and $i^{*}g$ commute with respect to the induced Poisson bracket. By definition of the inclusion $j$ and by (6.6), there exists $c\in{\mathbb{R}}$ such that
$T^{2}_{0}(i^{*}g)(z)=a_{x}{\mathfrak{R}}(z_{0}z_{1})+a_{y}{\mathfrak{I}}(z_{0}z_{1})+a_{\rho}(|z_{0}|^{2}+|z_{1}|^{2})+c\,i^{*}\Phi_{1}(z)$
for all $z=(z_{0},z_{1})\in{\mathbb{C}}^{2}$. Since $a_{x},a_{y},$ and $a_{\rho}$ are not all zero, the quadratic forms $i^{*}\Phi_{1}$ and $T^{2}_{0}(i^{*}g)$ are linearly independent. Hence the map $(i^{*}\Phi_{1},T^{2}_{0}(i^{*}g))\colon{\mathbb{C}}^{2}\to{\mathbb{R}}^{2}$ is regular on a dense subset. Since the Taylor polynomial of $i^{*}g-T^{2}_{0}(i^{*}g)$ of degree two at zero vanishes, the 2-form $d(i^{*}\Phi_{1})\wedge d(i^{*}g-T^{2}_{0}(i^{*}g))$ goes to 0 more quickly than $d(i^{*}\Phi_{1})\wedge d(T^{2}_{0}(i^{*}g))$ as $z=(z_{0},z_{1})$ approaches 0. Hence, the map $(i^{*}\Phi_{1},i^{*}g)$ is regular on a dense subset of an open neighborhood $W$ of 0 and so $(W,i^{*}\omega_{Y},{\mathbb{R}}\times{\mathbb{R}},(i^{*}\Phi_{1},i^{*}g))$ is an integrable system. Moreover, since $a_{x}^{2}+a_{y}^{2}-4a_{\rho}^{2}>0$, the point $0\in{\mathbb{C}}^{2}$ is a non-degenerate singular point with one focus-focus block. So by Lemma 5.4, there exist an open neighborhood $V$ of $0]\in{\mathbb{C}}^{2}/\!/S^{1}$ and a map $\psi\colon V\to{\mathbb{R}}^{2}$ taking $[0]$ to $(0,0)$ satisfying (2a) – (2c). Moreover, by Lemma 2.8, the inclusion $i$ induces a diffeomorphism $\overline{i}\colon{\mathbb{C}}^{2}/\!/S^{1}\to{M/\!/T}$. Setting $U:=\overline{i}(V)$ and $\Psi\colon U\to{\mathbb{R}}^{2}$ to be $\psi\circ\overline{i}^{-1}$, the result follows.
Case 4: $\mathbf{N>2}$. In this case, since the Taylor polynomial $T^{2}_{0}(j^{*}g)$ is $H$-invariant, by Remark 2.12 it lies in the span of $\{|z_{k}|^{2}\colon k=0,\dots,h\}$, as does $\Phi_{H}^{\xi}$ for any $\xi\in{\mathfrak{h}}$. Hence, any element in the span of $\{T^{2}_{0}(j^{*}g),\Phi_{H}^{\xi}\mid\xi\in{\mathfrak{h}}\}$ can be written as
$\sum\limits_{k=0}^{h}c_{k}|z_{k}|^{2}.$
The eigenvalues of the corresponding symplectic endomorphism are given by
$\pm 2c_{0}\sqrt{-1},\,\ldots,\,\pm 2c_{h}\sqrt{-1}.$
Since $p$ is non-degenerate, the span of $\{T^{2}_{0}(j^{*}g),\Phi_{H}^{\xi}\mid\xi\in{\mathfrak{h}}\}$ equals the span of $\{|z_{k}|^{2}\colon k=0,\dots,h\}$. Hence, the zero set of the map ${Y/\!/T}\to{\mathbb{R}}$ taking $t,\alpha,z]$ to $T^{2}_{0}(j^{*}g)$ is the point $[t,0,0]$. Moreover, since all the eigenvalues of the symplectic endomorphism corresponding to any element of the above span are imaginary, $p$ has purely elliptic type by Remark 6.1. ∎
A careful reading of the proof of Proposition 6.3 shows that it also demonstrates the following result, which we will need in a future paper.
Lemma 6.7.
Let $(Y,\omega_{Y},{\mathfrak{t}}\times{\mathbb{R}},f:=(\Phi_{Y},g))$ be an integrable system, where $Y=T\times_{H}{\mathfrak{h}}^{\circ}\times{\mathbb{C}}^{h+1}$ is a tall local model. Assume that $p=[1,0,0]$ is a non-degenerate critical point of $g$ modulo $\Phi$ with $g(p)=0$. If $j\colon{\mathbb{C}}^{h+1}\to Y$ is the inclusion $z\mapsto[1,0,z]$, then
  • the map ${Y/\!/T}\to{\mathbb{R}}$ taking $t,\alpha,z]\in{Y/\!/T}$ to $T^{1}_{0}(j^{*}g)(z)$ is identically zero, and
  • the zero set of the map ${Y/\!/T}\to{\mathbb{R}}$ taking $t,\alpha,z]\in{Y/\!/T}$ to $T^{2}_{0}(j^{*}g)(z)$ is
    • a single point if $p$ has purely elliptic type,
    • homeomorphic to $\{(x,y)\in{\mathbb{R}}^{2}\mid xy=0\}$ if $p$ has a hyperbolic block and connected $T$-stabilizer, and
    • homeomorphic to $\mathbb{R}$} otherwise.
7. Proof of the main result
In this section, we prove of the main results, Theorems 1.6 and 1.10. To begin, we explain how certain questions about the topology of the integrable systems we consider can be reduced to understanding Morse theory for oriented surfaces.
Proposition 7.1.
Let $(M,\omega,{\mathfrak{t}}\times{\mathbb{R}},f=(\Phi,g))$ be an integrable system such that $(M,\omega,\Phi)$ is a complexity one $T$-space. Assume that each tall singular point in $(M,\omega,{\mathfrak{t}}\times{\mathbb{R}},f)$ is non-degenerate. Fix $\beta\in{\mathfrak{t}}^{*}$. Each component of the reduced space $\Phi^{-1}(\beta)/T$ containing more than one point can be given the structure of a smooth oriented 2-dimensional manifold such that $\overline{g}\colon\Phi^{-1}(\beta)/T\to{\mathbb{R}}$ is a Morse function. Moreover, given $p\in\Phi^{-1}(\beta)$:
  1. (1)
    $[p]$ is a critical point of $\overline{g}$ of index $1$ exactly if $p$ has a hyperbolic block and connected $T$-stabilizer,
  2. (2)
    $[p]$ is a critical point of $\overline{g}$ of index 0 or $2$ exactly if $p$ is a critical point of $g$ modulo $\Phi$ with purely elliptic type, and
  3. (3)
    $[p]$ is a regular point of $\overline{g}$ exactly if $p$ has a focus-focus block, has a hyperbolic block with disconnected $T$-stabilizer, or is a regular point of $g$ modulo $\Phi$.
Proof.
We may assume that $\beta=0$. For simplicity, assume that $\Phi^{-1}(0)/T$ is connected and contains more than one point. Then, by definition, every point in $\Phi^{-1}(0)$ is tall. If $p\in\Phi^{-1}(0)$ is a critical point of $g$ modulo $\Phi$, then $p$ is also a singular point of $\left(M,\omega,{\mathfrak{t}}\times{\mathbb{R}},f\right)$. So, by assumption, $p$ is non-degenerate. Therefore, by Proposition 3.4, the claim follows from Propositions 4.5 and 6.3. ∎
Therefore, we review a few lemmas from Morse theory.
Lemma 7.2.
Let $N$ be a connected manifold and let $\overline{g}\colon N\to{\mathbb{R}}$ be a proper Morse function. If the fiber $\overline{g}^{-1}(c)$ is connected for each $c\in{\mathbb{R}}$, then $\overline{g}$ has at most one point of index 0 (or coindex 0).
Proof.
Let $p$ be a critical point of index 0. We may assume that $\overline{g}(p)=0$. By the Morse lemma, there exists a neighborhood $U$ of $p$ such that $\overline{g}^{-1}(0)\cap U=\{p\}$ and $\overline{g}(U)\subseteq[0,\infty)$. Since the fibers are connected, the former implies that $\overline{g}^{-1}(0)=\{p\}$. Hence, the latter implies that $\overline{g}^{-1}[0,\infty)$ is a nonempty subset of $N$ that is both open and closed. Since $N$ is connected, this implies that $\overline{g}$ achieves its minimum at $p$. Hence, $p$ is the unique point of index 0. ∎
Lemma 7.3.
Let $N$ be a closed, connected 2-dimensional manifold and let $\overline{g}\colon N\to{\mathbb{R}}$ be a Morse function. The following are equivalent:
  1. (1)
    no critical point of $\overline{g}$ has index one; and
  2. (2)
    the fiber $\overline{g}^{-1}(c)$ is connected for all $c\in{\mathbb{R}}$ and $N$ is simply connected.
Proof.
Assume that no critical point of $\overline{g}$ has index one. Since $N$ is 2-dimensional, by Proposition 5.6 the fiber $\overline{g}^{-1}(c)$ is connected for all $c\in{\mathbb{R}}$. Hence, since $N$ is closed, $\overline{g}$ has exactly two critical points by Lemma 7.2. By [Mil, Theorem 4.1], $N$ is homeomorphic to a sphere and, therefore, simply connected.
Conversely, assume that the fiber $\overline{g}^{-1}(c)$ is connected for all $c\in{\mathbb{R}}$ and that $N$ is simply connected. Since $N$ is closed, $\overline{g}$ has exactly one point of index zero and one point of index two by Lemma 7.2. Hence, by [Mil, Theorem 5.2], the Euler characteristic of $N$ is equal to $2-C_{1}$, where $C_{1}$ is the number of critical points of $\overline{g}$ of index one. Since $N$ is closed and simply connected, it follows that $C_{1}=0$. ∎
We can now prove our main result.
Proof of Theorem 1.6.
Let $\left(M,\omega,{\mathfrak{t}}\times{\mathbb{R}},f:=(\Phi,g)\right)$ be an integrable system such that $(M,\omega,\Phi)$ is a complexity one $T$-space with a proper moment map. Assume that each tall singular point in $\left(M,\omega,{\mathfrak{t}}\times{\mathbb{R}},f\right)$ is non-degenerate and that no such point has a hyperbolic block and connected $T$-stabilizer. Fix $\beta\in{\mathfrak{t}}^{*}$ such that $\Phi^{-1}(\beta)/T$ contains more than one point; we may assume that $\beta=0$. Since $\Phi$ is proper, the reduced space M//T{M/\!/T} is compact and connected by Theorem 5.5. Hence, every point in M//T{M/\!/T} is tall, and so by Proposition 7.1, M//T{M/\!/T} is a smooth closed, connected oriented surface so that $\overline{g}\colon{M/\!/T}\to{\mathbb{R}}$ is a Morse function; moreover, $\overline{g}$ has no critical points of index $1$. Therefore, by Lemma 7.3 the fiber $\overline{g}^{-1}(c)$ is connected for all $c\in{\mathbb{R}}$ and M//T{M/\!/T} is simply connected. Finally, since $T$ is connected, this implies that the fiber $f^{-1}(0,c)$ is connected for all $c\in{\mathbb{R}}$. ∎
To prove Theorem 1.10, we also need the following result.
Lemma 7.4.
Let $N$ be a closed, connected oriented $2$-dimensional manifold and let $\overline{g}\colon N\to{\mathbb{R}}$ be a Morse function. If the fiber $\overline{g}^{-1}(c)$ is connected for each $c\in{\mathbb{R}}$, then each fiber contains an even number of critical points of index $1$.
Proof.
By Lemma 7.2, $N$ has a unique critical point of index 0 (at the minimum), and a unique critical point of index $2$ (at the maximum). Assume that the fiber $\overline{g}^{-1}(c)$ contains $k\geq 1$ critical points of index $1$. By the Morse Lemma $c$ is in the interior of $\overline{g}(N)$, and so there are no critical points of index 0 or index $2$ in $\overline{g}^{-1}(c)$. Fix $\epsilon>0$ so that $c-\epsilon,c+\epsilon]$ contains no other critical values. Since there is only one point of index 0 we know that the preimage $\overline{g}^{-1}((-\infty,c-\epsilon])$ is connected. Hence, the first Betti number of $\overline{g}^{-1}((-\infty,c+\epsilon])$ is the first Betti number of $\overline{g}^{-1}((-\infty,c-\epsilon])$ plus $k$. Since every compact oriented $2$-dimensional manifold with boundary with an odd first Betti number has at least two boundary components, both of these Betti number must be even. Hence $k$ is even. ∎
Proof of Theorem 1.10.
Let $\left(M,\omega,{\mathfrak{t}}\times{\mathbb{R}},f:=(\Phi,g)\right)$ be an integrable system such that $(M,\omega,\Phi)$ is a complexity one $T$-space with a proper moment map. Assume that each tall singular point in $\left(M,\omega,{\mathfrak{t}}\times{\mathbb{R}},f\right)$ is non-degenerate. Fix $\beta\in{\mathfrak{t}}^{*}$ such that $\Phi^{-1}(\beta)/T$ contains more than one point; we may assume that $\beta=0$. Since $\Phi$ is proper, the reduced space M//T{M/\!/T} is compact and connected by Theorem 5.5. Hence, every point in M//T{M/\!/T} is tall, and so by Proposition 7.1, M//T{M/\!/T} can be given the structure of a smooth closed, connected oriented surface so that $\overline{g}\colon{M/\!/T}\to{\mathbb{R}}$ is a Morse function with connected fibers; moreover, given $p\in\Phi^{-1}(0)$, the orbit $[p]$ is a critical point of $\overline{g}$ of index $1$ exactly if $p$ has a hyperbolic block and connected $T$-stabilizer. Finally, since $T$ is connected, the fiber $\overline{g}^{-1}(c)$ is connected exactly if the fiber $f^{-1}(\beta,c)$ is connected for any $c\in{\mathbb{R}}$.
If $\Phi^{-1}(\beta_{0})/T$ is simply connected for some $\beta_{0}\in\Phi(M)$, then M//T{M/\!/T} is simply connected by [KT1, Lemma 5.7 and Corollary 9.7]. Therefore, the fibers $f^{-1}(\beta,c)$ are connected for all $c\in{\mathbb{R}}$ exactly if M//T{M/\!/T} has no tall singular points with a hyperbolic block and connected $T$-stabilizer by Lemma 7.3.
Similarly, if each fiber of $f$ contains at most one $T$-orbit of non-degenerate singular points with a hyperbolic block, then each fiber of $\overline{g}$ contains at most one one critical point of index $1$. Therefore, the fibers $f^{-1}(\beta,c)$ are connected for all $c\in{\mathbb{R}}$ exactly if M//T{M/\!/T} has no tall singular points with a hyperbolic block and connected $T$-stabilizer by Lemmas 7.3 and 7.4. ∎
Since Proposition 5.6 is of central importance, we conclude this section with a slight variation of the proof given in [PRV]. We begin with a preliminary lemma (cf. [PRV, Lemma 4.3]).
Lemma 7.5.
Let $N$ be a connected manifold and let $h\colon N\to{\mathbb{R}}$ be a proper function such that $h(N)\subseteq[0,\infty)$. Assume that the restriction of $h$ to $h^{-1}(0,\infty)$ is a Morse-Bott function with no points of index $1$, and that 0 is an isolated critical value. Then $h^{-1}(0)$ is connected.
Proof.
Consider points $x$ and $y$ in $h^{-1}(0)$. Since $N$ is path-connected, there is a path $\gamma$ from $x$ to $y$. Hence, $x$ and $y$ are in the same component of $h^{-1}([0,b))$ for some $b\in{\mathbb{R}}$. If $C\subseteq h^{-1}((0,\infty))$ is a critical submanifold, then since the index of $C$ is not $1$, the restriction $H^{0}(h^{-1}([0,f(C)+\epsilon)))\to H^{0}(h^{-1}([0,f(C)-\epsilon)))$ is a surjection for sufficiently small $\epsilon>0$. In particular, if $x$ and $y$ are in the same component of $h^{-1}([0,h(C)+\epsilon))$, then they are also in the same component of $h^{-1}([0,h(C)-\epsilon))$. Moreover, since $h$ is proper and 0 is an isolated critical value, there are a finite number of such critical submanifolds in $h^{-1}([0,b))$. Hence, by induction, $x$ and $y$ are in the same component of $h^{-1}(0).$
Proof of Proposition 5.6.
Fix $c\in{\mathbb{R}}$ and define $h\colon N\to{\mathbb{R}}$ by $h=(\overline{g}-c)^{2}$. On the set $\overline{g}^{-1}(c,\infty)$ the function $h$ is Morse-Bott; the critical components of $h$ and $\overline{g}$ coincide and have the same index. On the set $\overline{g}^{-1}(-\infty,c)$ the function $h$ is Morse-Bott; the critical components of $h$ and $\overline{g}$ coincide but the index and coindex of each critical manifold are exchanged. Finally, since $h$ is a proper Morse-Bott function there exists a neighborhood of $c$ that contains no other critical values, so 0 is an isolated critical value. Therefore, $h^{-1}(0)=\overline{g}^{-1}(c)$ is connected by Lemma 7.5. ∎